### v1

Abstract
This paper studies the linear quadratic regulation (LQR) problem of unknown discrete-time systems via dynamic output feedback learning control. In contrast to the state feedback, the optimality of the dynamic output feedback control for solving the LQR problem requires an implicit condition on the convergence of the state observer. Moreover, due to unknown system matrices and the existence of observer error, it is difficult to analyze the convergence and stability of most existing output feedback learning-based control methods. To tackle these issues, we propose a generalized dynamic output feedback learning control approach with guaranteed convergence, stability, and optimality performance for solving the LQR problem of unknown discrete-time linear systems. In particular, a dynamic output feedback controller is designed to be equivalent to a state feedback controller. This equivalence relationship is an inherent property without requiring convergence of the estimated state by the state observer, which plays a key role in establishing the off-policy learning control approaches. By value iteration and policy iteration schemes, the adaptive dynamic programming based learning control approaches are developed to estimate the optimal feedback control gain. In addition, a model-free stability criterion is provided by finding a nonsingular parameterization matrix, which contributes to establishing a switched iteration scheme. Furthermore, the convergence, stability, and optimality analyses of the proposed output feedback learning control approaches are given. Finally, the theoretical results are validated by two numerical examples.
keywords:
Output feedback, linear quadratic regulation, policy iteration, value iteration, adaptive dynamic programming.
journal: arXiv
1 Introduction
Reinforcement learning (RL) (Sutton & Barto, 2018; Lewis et al., 2012a), also known as adaptive dynamic programming (ADP) (Jiang & Jiang, 2017; Powell, 2004; Liu et al., 2021; Lewis et al., 2012b) in the field of control, is one of the most important cutting-edge research fields in artificial intelligence for optimization Wu et al. (2025) and decision-making in uncertain environments. Especially, machine learning method has shown promise in monitoring real-time green carbon dioxide tracking in refinery processes Cao et al. (2025). Moreover, the ADP algorithm offers iteration learning frameworks to estimate the optimal control policy for solving several optimal control problems by only accessing a series of historical data of input, output, and/or state instead of prior knowledge of system dynamics; see, Vrabie et al. (2009); Jiang & Jiang (2012); Modares et al. (2016); Chen et al. (2023); Modares & Lewis (2014); Gao & Jiang (2016); Yang et al. (2018); Rizvi & Lin (2020); Duan et al. (2024), and references therein. As an ongoing research topic, it is of great significance to advance the development of the ADP method with output feedback control, as it has the potential to effectively address the optimal control problem under the unmeasurable state in various practical settings. However, it is shown in Duan et al. (2024); Rizvi & Lin (2023); Chen et al. (2023) that the theoretical guarantees of the ADP-based output feedback learning control methods, including convergence, stability, and optimality, remain substantial challenges.
As is widely known, the linear quadratic regulator is a full-state feedback controller, and its optimal control gain can be obtained by solving a corresponding algebraic Riccati equation (ARE) (Kalman et al., 1960; Lancaster & Rodman, 1995) which relies on accurate knowledge of the system matrices. For the purpose of obtaining the optimal state feedback control policy without identifying system matrices, a novel ADP-based learning algorithm was first proposed in Vrabie et al. (2009) to solve the LQR problem of partially unknown linear systems. Using a single state trajectory, a policy iteration (PI) scheme was employed in Vrabie et al. (2009) to estimate the optimal control policy in an online fashion, and the convergence is guaranteed by requiring an initial stabilizing control policy and the persistence of excitation (PE) condition. Later on, by using an integral RL technique with the available data information of state and input, an ADP-based off-policy model-free learning algorithm with the PI scheme was proposed in Jiang & Jiang (2012) to approximate the optimal state feedback controller for solving the LQR problem of completely unknown linear systems. To relax the requirement of initial admissible control policy, a value iteration (VI) based learning algorithm was proposed in Bian & Jiang (2016) at the cost of convergence rate. In addition, the stability analysis of the ADP-based state feedback learning control approach was provided in Bian & Jiang (2016). It is worth noting that both the convergence and optimality of the proposed learning algorithms in Jiang & Jiang (2012) and Bian & Jiang (2016) can be guaranteed under a rank condition which is a milder condition than the PE condition. The ADP-based learning framework in Vrabie et al. (2009); Jiang & Jiang (2012); Bian & Jiang (2016) for solving the LQR problem has been extended to solve several optimal control problems, for instance, the optimal tracking problem (Modares & Lewis, 2014), the optimal robust control problem (Pang et al., 2022), and the optimal output regulation problem (Gao & Jiang, 2016; Odekunle et al., 2020). These ADP-based state feedback learning control approaches provide the fundamental results for developing the output feedback results.
Based on the existing results on model-free state feedback (Vrabie et al., 2009; Jiang & Jiang, 2012; Bian & Jiang, 2016), several output feedback learning control approaches were proposed by designing a stable state observer with parameterization methods, for instance, Lewis & Vamvoudakis (2011); Modares et al. (2016); Chen et al. (2022, 2023); Rizvi & Lin (2020, 2023); Gao & Jiang (2019); Gao et al. (2016). In particular, a state parameterization method for linear discrete-time systems was first designed in Lewis & Vamvoudakis (2011), which makes the output feedback controller with the constructed state equivalent to a state feedback controller. Similar to the state feedback learning control in Vrabie et al. (2009), by using the measurement of the output and input data with a probing noise to replace the system matrices in the corresponding ARE, a model-free learning control approach was proposed to solve the LQR problem in Lewis & Vamvoudakis (2011). It should be pointed out that the probing noise would lead to an estimation error in the learning process. The framework of output feedback control with the state parameterization method and the ADP algorithm was widely extended to solve several optimal control problems, for example, the optimal robust control problem (Gao et al., 2016) and the optimal output regulation problem (Gao & Jiang, 2019). Nevertheless, as stated in Postoyan et al. (2016), the discounted controller in Lewis & Vamvoudakis (2011); Modares et al. (2016) is not an optimal solution and may not even be a stabilizing one if the discounting factor in the cost function is not carefully chosen.
Different from Lewis & Vamvoudakis (2011); Modares et al. (2016), a novel state parameterization method was developed in Rizvi & Lin (2019) to solve the output feedback LQR problem of discrete-time linear systems. In particular, an internal system was embedded in the controller that can be seen as a stable state observer under the observability condition. In addition, the problem of exploration noise bias was avoided in Rizvi & Lin (2019) without resorting to a discounting factor by proposing a model-free Q-learning output feedback control method with both PI and VI schemes. This output feedback learning control framework designed in Rizvi & Lin (2019) has been expanded to solve the tracking problems (Chen et al., 2022, 2023) and the zero-sum games (Rizvi & Lin, 2020). However, as stated in Rizvi & Lin (2023), the convergence and optimality of the output feedback learning control algorithm in the aforementioned references (Rizvi & Lin, 2019, 2020; Chen et al., 2022, 2023) may not be guaranteed unless the convergence of the state observer is achieved. Moreover, stability analysis has not been provided in these works, while it is a bare requirement in a closed-loop control design.
In this paper, we aim to give a generalized result of the ADP-based output feedback learning control for solving the discrete-time LQR problem. The main significance of the result is that we provide an optimal solution to the output feedback LQR problem with convergence guaranteed, and the stability analysis is also given for the learning process. Compared to the recent works on ADP-based output feedback learning control (Rizvi & Lin, 2019, 2023; Chen et al., 2022), the contributions of this paper are threefold.
  • 1)
    A generalized dynamic output feedback learning controller is proposed, which has an inherent property of being equivalent to a state feedback controller. The equivalence relationship holds without the requirement of the convergence of state observer, which not only contributes to establishing the ADP-based learning algorithms but also ensures the optimality of the proposed dynamic output feedback learning controller.
  • 2)
    Two model-free ADP-based learning algorithms are proposed by employing the PI scheme and the VI scheme, respectively. The convergence of the proposed learning algorithms with both PI and VI can be guaranteed under the rank conditions, and it is immune to the exploration noise and the state observer error. In addition, the rank condition of state parameterization can be achieved with more flexible pole assignment configurations.
  • 3)
    A model-free stability criterion is provided, and the stability analysis of the learning algorithm is given. The stability criterion paves the way for us to propose a switched iteration (SI) scheme integrated with PI and VI to solve the LQR problem. The proposed SI scheme can relax the requirement of an initial admissible controller in the PI scheme and accelerate the convergence rate in the VI scheme.
The rest of this paper is organized as follows. Section 2 presents the preliminaries and problem formulation. In Section 3, a novel dynamic output feedback control design method is proposed, and the ADP-based learning control approach is developed. The stability, convergence, and optimality analysis are also presented in this section. Finally, the simulation results are given in Section 5, and Section 6 draws the conclusion.
2 Preliminaries and Problem Formulation
2.1 Notation
[…]
2.2 Plant and Preliminaries
Consider the following linear discrete-time systems
$x({k+1})$ $=Ax(k)+Bu(k)$ (1a)
$y(k)$ $=Cx(k)$ (1b)
where $x\in\mathbb{R}^{n}$, $u\in\mathbb{R}^{m}$, and $y\in\mathbb{R}^{p}$ are the unmeasurable state, the control input, and the output, respectively. $A$, $B$, and $C$ are constant system matrices with appropriate dimensions and satisfy the following condition.
Assumption 1
The pair $(A,B)$ is stabilizable, and the pair $(A,C)$ is observable.
Define the utility function as
$\mathcal{C}_{y}(y(k),u(k))=y^{\mbox{\tiny T}}(k){Q_{y}}y(k)+u^{\mbox{\tiny T}}(k){R}u(k)$ (2)
where $Q_{y}=Q_{y}^{\mbox{\tiny T}}\in\mathbb{R}^{p\times p}\geq\mathbf{0}$, $R=R^{\mbox{\tiny T}}\in\mathbb{R}^{m\times m}>\mathbf{0}$ are the weight matrices, and the pair $(A,\sqrt{C^{\mbox{\tiny T}}Q_{y}C})$ is observable. The objective of the LQR problem is to find a feedback control sequence $u(k)$ such that the following infinite horizon quadratic value function
$V(x(0))=\sum_{k=0}^{\infty}y^{\mbox{\tiny T}}(k){Q_{y}}y(k)+u^{\mbox{\tiny T}}(k){R}u(k)$ (3)
is minimized.
If the state $x(k)$ is measurable, the LQR problem can be solved by the well-known full-state feedback controller
$u(k)=-{K^{*}}{x(k)}$ (4)
where the optimal control gain is
$K^{*}=\left(R+B^{\mbox{\tiny T}}P^{*}B\right)^{-1}B^{\mbox{\tiny T}}P^{*}A$ (5)
with the symmetric matrix $P^{*}$ as the solution to the discrete algebraic Riccati equation (ARE) (Lancaster & Rodman, 1995)
${A}^{\mbox{\tiny T}}{P}A-P+C^{\mbox{\tiny T}}Q_{y}C-A^{\mbox{\tiny T}}PB\left(R+B^{\mbox{\tiny T}}PB\right)^{-1}B^{\mbox{\tiny T}}PA=\mathbf{0}.$ (6)
Then, the value function (3) under the optimal controller (4) is given as
$V(x(0))=x^{\mbox{\tiny T}}(0)P^{*}x(0).$ (7)
That is, the LQR problem can be solved by computing ARE (6) and (5) when the state is available. Based on this result, it is not difficult to know that the LQR problem with unmeasurable state can be solved by reconstructing or estimating the real state according to the separation principle (Lewis et al., 2012a). As given in Lewis & Vamvoudakis (2011, Lemma 1), under the observability condition, the state of system (1) can be reconstructed by
$x(k)=$ $M_{\bar{y}}\bar{y}_{k-1,k-N}+M_{\bar{u}}\bar{u}_{k-1,k-N}=\left[M_{\bar{y}},~{}M_{\bar{u}}\right]\left[\begin{matrix}\bar{y}_{k-1,k-N}\\ \bar{u}_{k-1,k-N}\end{matrix}\right]$
::=:= $\bar{M}\bar{\upsilon}_{k-1,k-N}$ (8)
where $N$ is a positive integer which is greater than the observability index of system (1), $M_{\bar{y}}=A^{N}\left(V_{N}^{\mbox{\tiny T}}V_{N}\right)^{-1}V_{N}^{\mbox{\tiny T}}$, $M_{\bar{u}}=U_{N}-M_{\bar{y}}T_{N}$, and
$\bar{y}_{k-1,k-N}=$ ${\mbox{col}}\left(y(k-1),\dots,y(k-N-1),y(k-N)\right)$
$\bar{u}_{k-1,k-N}=$ ${\mbox{col}}\left(u(k-1),\dots,u(k-N-1),u(k-N)\right)$
with
$U_{N}=$ $\left[B,~{}AB,~{}\cdots,~{}A^{N-1}B\right]$
$V_{N}=$ $\left[\begin{matrix}CA^{N-1}\\ \vdots\\ CA\\ C\end{matrix}\right],~{}~{}T_{N}=\left[\begin{matrix}\mathbf{0}&CB&CAB&\cdots&CA^{N-2}B\\ \mathbf{0}&\mathbf{0}&CB&\cdots&CA^{N-3}B\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\cdots&CB\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\cdots&\mathbf{0}\\ \end{matrix}\right].$
Using the state reconstruction in (2.2), the LQR problem can be solved by the following optimal controller:
$u(k)=-K^{*}\bar{M}\bar{\upsilon}_{k-1,k-N}:=-K_{\bar{M}}\bar{\upsilon}_{k-1,k-N}.$ (9)
Different from the state reconstruction given in (2.2) which is specific to the discrete-time systems, an alternative way to obtain the state information is to use a state observer to estimate the real state. Concretely, the state of system (1) can be estimated by using the following Luenburger observer:
$\hat{x}(k+1)=(A-LC)\hat{x}(k)+Bu(k)+Ly(k).$ (10)
It is known that the Luenburger observer (10) has the property that $\lim_{k\rightarrow\infty}\hat{x}(k)=x(k)$ when all the eigenvalues of $(A-LC)$ are set within the unit circle under the observability condition. As given in Rizvi & Lin (2019), the estimated state $\hat{x}(k)$ can be parametrized by
$\hat{x}(k)=W_{y}\omega(k)+W_{u}\sigma(k)=\left[W_{y},~{}W_{u}\right]\left[\begin{matrix}\omega(k)\\ \sigma(k)\end{matrix}\right]:=\bar{W}\varpi(k)$ (11)
where the parameterization matrices $W_{y}\in\mathbb{R}^{n\times np}$ and $W_{u}\in\mathbb{R}^{n\times nm}$ are determined by the coefficients of the numerators in the transfer function matrix of the Luenburger observer system (10) with $y(k)$ and $u(k)$ being the inputs to the observer system, and the dynamics of $\omega(k)$ and $\sigma(k)$ are given as
$\omega(k+1)=$ $~{}{\bar{\mathcal{A}}}_{y}\omega(k)+{\bar{\mathcal{B}}}_{y}y(k),~{}\omega(0)=\mathbf{0}$
$\sigma(k+1)=$ $~{}{\bar{\mathcal{A}}}_{u}\sigma(k)+{\bar{\mathcal{B}}}_{u}u(k),~{}\sigma(0)=\mathbf{0}$
where ${\bar{\mathcal{A}}}_{y}\in\mathbb{R}^{np\times np}$ and ${\bar{\mathcal{A}}}_{u}\in\mathbb{R}^{nm\times nm}$ are user-defined matrices determined by the eigenvalues of $(A-LC)$, and ¯${\bar{\mathcal{B}}}_{y}\in\mathbb{R}^{np\times p}$ and ¯${\bar{\mathcal{B}}}_{u}\in\mathbb{R}^{nm\times m}$ are also user-defined matrices which can make $({\bar{\mathcal{A}}}_{y},~{}{\bar{\mathcal{B}}}_{y})$ and $({\bar{\mathcal{A}}}_{u},~{}{\bar{\mathcal{B}}}_{u})$ controllable. It is easy to see that the data of $\omega(k)$ and $\sigma(k)$ are available since their system matrices $({\bar{\mathcal{A}}}_{y},~{}{\bar{\mathcal{B}}}_{y})$ and $({\bar{\mathcal{A}}}_{u},~{}{\bar{\mathcal{B}}}_{u})$ are user-defined.
Using the state parameterization in (11), the LQR problem can be solved by
$u(k)=-K^{*}\bar{W}\varpi(k):=-K_{\bar{W}}\varpi(k)$ (12)
when the state of the observer system converges to the real state. It is worth mentioning that, if all the eigenvalues of $(A-LC)$ are chosen as 00, the observer state $\hat{x}(k)$ converges to the real state $x(k)$ when $k\geq n$.
Thus, by using the state reconstruction in (2.2) or the state parameterization in (11), the LQR problem can be solved by the output feedback controller (9) or (12) when the accurate knowledge of system matrices are known.
2.3 Problem Formulation
In this paper, we aim to solve the LQR problem of linear discrete-time systems with completely unknown system matrices and unmeasurable state, which can be formulated as follows:
Problem 1
For system (1) where the system matrices $A$, $B$, and $C$ are unknown, and the state $x$ is unmeasurable, find an optimal control policy sequence $u(k)$ to satisfy
$V(x(0))=\sum_{k=0}^{\infty}y^{\mbox{\tiny T}}(k){Q_{y}}y(k)+u^{\mbox{\tiny T}}(k){R}u(k)=x^{\mbox{\tiny T}}(0)P^{*}x(0)$
where $P^{*}$ is the solution to ARE (6).
The challenges of Problem 1 arise from the solution of ARE (6) and the design of the parameterization matrix $\bar{M}$ or $\bar{W}$, which becomes difficult in the absence of knowledge of the system matrices $A$, $B$, and $C$.
By combining the ADP method with a series of historical data of input and output, some ADP-based output feedback learning approaches focused on estimating the optimal control gain $K_{\bar{M}}$ and $K_{\bar{W}}$ directly, for instance, Lewis & Vamvoudakis (2011); Rizvi & Lin (2019, 2020, 2023); Gao et al. (2016); Chen et al. (2023). Particularly, it is proven in Rizvi & Lin (2023) that the requirement on the full row rank of $\bar{W}$ is necessary to guarantee the convergence performance of the aforementioned ADP-based output feedback learning approaches, so is the same requirement on $\bar{M}$ along the similar proof. However, as we can see from (2.2), under Assumption 1, the parameterization matrix $\bar{M}$ may not be of full row rank when the system matrix $A$ has the eigenvalue 00. It follows from Rizvi & Lin (2023, Theorem 4) that the parameterization matrix $\bar{W}$ may not be of full row rank when the matrices $(A-LC)$ and $A$ have common eigenvalues. On the other hand, as stated in Postoyan et al. (2016), the stability of the closed-loop system under the learning control approach proposed in Lewis & Vamvoudakis (2011) cannot be guaranteed. Although this issue was eliminated in Rizvi & Lin (2019, 2023); Chen et al. (2023), the requirement of the convergence of state observer is needed to ensure the equivalence of the dynamic output feedback controller with the state feedback controller. This implies that the observer error will influence the convergence and optimality performance of the ADP-based output feedback learning control approaches in Rizvi & Lin (2019, 2023); Chen et al. (2023).
To deal with the above issues, the objective of this paper is to provide a generalized learning based output feedback solution to Problem 1 and a detailed analysis of convergence, stability, and optimality for the proposed output feedback learning control approach.
3 Main Results
In this section, we propose a novel output feedback learning control approach to solve the LQR problem. In particular, a new dynamic output feedback controller is designed that is equal to a static state feedback controller. Then, a data-driven learning algorithm is established to estimate the optimal control gain without prior knowledge of system matrices. Finally, the convergence, stability, and optimality analyses of the proposed learning control approach are given.
3.1 Dynamic Output Feedback Controller Design
Now we present a generalized dynamic output feedback controller design method, where the internal model is established without using any prior knowledge of the system dynamics. Moreover, the equivalence relationship between the proposed output feedback controller and the state feedback controller always holds even in the presence of observer error.
Theorem 1
Under Assumption 1, there exists a dynamic output feedback controller
$u(k)$ $=-\mathcal{K}\eta(k)$ (13a)
$\eta(k+1)$ $=\mathcal{G}_{1}\eta(k)+\mathcal{G}_{2}\zeta(k)$ (13b)
that is equal to a static state feedback controller $u(k)=-Kx(k)$, $\forall k=0,1,2,\dots$, where $\eta\in\mathbb{R}^{n(m+p+1)}$ is the state of the internal model (13b) with a user-defined matrix pair $(\mathcal{G}_{1},\mathcal{G}_{2})$, $\mathcal{K}=KM$ is the control gain with a parameterization constant matrix $M$, and $\zeta=\textnormal{{\mbox{col}}}(u,y,0)$.
Proof. Consider a Luenburger observer system described as in (10) and rewiritten as
$\hat{x}(k+1)=A_{L}\hat{x}(k)+Bu(k)+Ly(k),~{}~{}\hat{x}(0)=0$ (14)
where $A_{L}=A-LC$ is the observer system matrix with observer gain $L$. By Assumption 1, the characteristic polynomial of $A_{L}$ can be designed arbitrarily, i.e.,
$\Lambda(z)\!=\!\mbox{det}(zI_{n}\!-\!A_{L})\!=\!z^{n}\!+\!\alpha_{n\!-\!1}z^{n\!-\!1}\!+\!\cdots\!+\!\alpha_{1}z\!+\!\alpha_{0}$ (15)
is user-defined with known coefficients $\alpha_{0},\alpha_{1},\dots,\alpha_{n-1}$.
Define the error as
$\varepsilon_{x}(k)=x(k)-\hat{x}(k).$ (16)
If follows from (1) and (14) that the dynamics of error is
$\varepsilon_{x}(k+1)=A_{L}\varepsilon_{x}(k).$ (17)
Using the $z$-transformation, the explicit solution to $\varepsilon_{x}(k)$ in (17) is given as
$\varepsilon_{x}(k)$ $=\mathcal{Z}^{-1}\left(\frac{\varepsilon_{x}(0)}{(zI_{n}-A_{L})}\right)=\mathcal{Z}^{-1}\left(\frac{\mbox{adj}(zI_{n}-A_{L})\varepsilon_{x}(0)}{\mbox{det}(zI_{n}-A_{L})}\right)$
$=$ $\mathcal{Z}^{-1}\left(\left[\begin{matrix}\frac{a^{1}_{n-1}z^{n-1}+a^{1}_{n-2}z^{n-2}+\cdots+a^{1}_{1}z+a^{1}_{0}}{z^{n}+\alpha_{n-1}z^{n-1}+\cdots+\alpha_{1}z+\alpha_{0}}\\ \frac{a^{2}_{n-1}z^{n-1}+a^{2}_{n-2}z^{n-2}+\cdots+a^{2}_{1}z+a^{2}_{0}}{z^{n}+\alpha_{n-1}z^{n-1}+\cdots+\alpha_{1}z+\alpha_{0}}\\ \vdots\\ \frac{a^{n}_{n-1}z^{n-1}+a^{n}_{n-2}z^{n-2}+\cdots+a^{n}_{1}z+a^{n}_{0}}{z^{n}+\alpha_{n-1}z^{n-1}+\cdots+\alpha_{1}z+\alpha_{0}}\\ \end{matrix}\right]\right)$
$=$ $\mathcal{Z}^{-1}\left(\left[\begin{matrix}a^{1}_{0}&a^{1}_{1}&\cdots&a^{1}_{n-1}\\ a^{2}_{0}&a^{2}_{1}&\cdots&a^{2}_{n-1}\\ \vdots&\vdots&\cdots&\vdots\\ a^{n}_{0}&a^{n}_{1}&\cdots&a^{n}_{n-1}\\ \end{matrix}\right]\left[\begin{matrix}\frac{1}{\Lambda(z)}\\ \frac{z}{\Lambda(z)}\\ \vdots\\ \frac{z^{n-1}}{\Lambda(z)}\\ \end{matrix}\right]\right)$
::=:= $M_{\varepsilon_{x}}(A,L,C,\varepsilon_{0})\mathcal{Z}^{-1}\left(\left[\begin{matrix}\frac{1}{\Lambda(z)}\\ \frac{z}{\Lambda(z)}\\ \vdots\\ \frac{z^{n-1}}{\Lambda(z)}\\ \end{matrix}\right]\right)$ (18)
where $M_{\varepsilon_{x}}(A,L,C,\varepsilon_{x}(0))$ is an unknown constant matrix determined by $(A,L,C,\varepsilon_{x}(0))$.
Define a new system described by
$\eta_{\varepsilon}(k+1)=\mathcal{A}_{\varepsilon}\eta_{\varepsilon}(k)$
where the system matrix $\mathcal{A}_{\varepsilon}$ is known and has the same polynomial as $A_{L}$. Along the same derivation as in (3.1), the explicit solution to $\eta_{\varepsilon}(k)$ is
$\eta_{\varepsilon}(k)=M_{\eta_{\varepsilon}}(\mathcal{A}_{\varepsilon},\eta_{\varepsilon}(0))\mathcal{Z}^{-1}\left(\left[\begin{matrix}\frac{1}{\Lambda(z)}\\ \frac{z}{\Lambda(z)}\\ \vdots\\ \frac{z^{n-1}}{\Lambda(z)}\end{matrix}\right]\right)$
where $M_{\eta_{\varepsilon}}(\mathcal{A}_{\varepsilon},\eta_{\varepsilon}(0))$ is a computable constant matrix determined by the user-defined $\mathcal{A}_{\varepsilon}$ and $\eta_{\varepsilon}(0)$. It is easy to choose $\eta_{\varepsilon}(0)$ satisfying $M_{z_{\varepsilon}}(\mathcal{A}_{\varepsilon},\eta_{\varepsilon}(0))$ of full rank. Thus, we have
$\varepsilon_{x}(k)=M_{\varepsilon_{x}}M_{\eta_{\varepsilon}}^{-1}\eta_{\varepsilon}(k):=M_{\varepsilon_{x}\leftarrow\eta_{\varepsilon}}\eta_{\varepsilon}(k),~{}~{}~{}\forall k=0,1,2,\dots$ (19)
where $M_{\varepsilon_{x}\leftarrow\eta_{\varepsilon}}$ is an unknown parameterization transformation matrix.
Along the similar derivation in Rizvi & Lin (2023), using the $z$-transformation, the explicit solution to $\hat{x}(k)$ in (14) is given as
$\hat{x}(k)$ $=\mathcal{Z}^{-1}\left(\sum_{i=1}^{m}\frac{B_{i}U_{i}(z)}{(zI_{n}-A_{L})}+\sum_{j=1}^{p}\frac{L_{j}Y_{j}(z)}{(zI_{n}-A_{L})}\right)$
$=\sum_{i=1}^{m}M_{u}^{i}\eta_{u}^{i}(k)+\sum_{j=1}^{p}M_{y}^{j}\eta_{y}^{j}(k)$ (20)
where $M_{u}^{i}\in\mathbb{R}^{n\times{n}}$ and $M_{y}^{j}\in\mathbb{R}^{n\times{n}}$ are unknown constant matrices determined by the matrices $A$, $L$, $C$, and $B$, and $\eta_{u}^{i}(k)$ and $\eta_{y}^{j}(k)$ are the internal states with the dynamics as follows:
$\eta_{u}^{i}(k+1)$ $=\mathcal{A}z_{u}^{i}(k)+bu_{i}(k)$
$\eta_{y}^{j}(k+1)$ $=\mathcal{A}z_{y}^{j}(k)+by_{j}(k)$
with
$\mathcal{A}=\left[\!\!\begin{array}[]{ccccc}0&1&0&\cdots&0\\ 0&0&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&1\\ -\alpha_{0}&-\alpha_{1}&-\alpha_{2}&\cdots&-\alpha_{n-1}\end{array}\!\!\right],~{}b=\left[\!\!\begin{array}[]{c}0\\ 0\\ \vdots\\ 0\\ 1\end{array}\!\!\right]$ (31)
where $\mathcal{A}$ is determined by the coefficients of $\Lambda(s)$ as in (15).
Combining (16), (19), and (3.1), we have
$x(k)$ $=\hat{x}(k)+\varepsilon_{x}(k)$
$=\sum_{i=1}^{m}M_{u}^{i}\eta_{u}^{i}(k)+\sum_{j=1}^{p}M_{y}^{j}\eta_{y}^{j}(k)+M_{\varepsilon_{x}\leftarrow\eta_{\varepsilon}}\eta_{\varepsilon}(k)$
$:=M\eta(k),~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}k=0,1,\dots$ (32)
where
$M$ $\!=\![M_{u}^{1},\dots,M_{u}^{m},M_{y}^{1},\dots,M_{y}^{p},M_{\varepsilon_{x}\leftarrow\eta_{\varepsilon}}]\!\in\!\mathbb{R}^{n\times n_{z}}$ (33)
$\eta$ $={\mbox{col}}\left(\eta_{u}^{1},\dots,z_{u}^{m},z_{y}^{1},\dots,z_{y}^{p},\eta_{\varepsilon}\right)\in\mathbb{R}^{n_{z}}$ (34)
with $n_{z}=n(m+p+1)$.
Finally, the dynamic output feedback controller is designed as
$u(k)$ $=-\mathcal{K}\eta(k)$ (35a)
$\eta(k+1)$ $=\mathcal{G}_{1}\eta(k)+\mathcal{G}_{2}\zeta(k)$ (35b)
where $\mathcal{K}=KM$ is the control gain with $M$ defined in (33), $\zeta={\mbox{col}}(u,y,0)$, and
$\mathcal{G}_{1}=\mbox{block diag}[\underbrace{\mathcal{A},\cdots,\mathcal{A}}_{m+p},\mathcal{A}_{\varepsilon}],~{}\mathcal{G}_{2}=\mbox{block diag}[\underbrace{b,\cdots,b}_{m+p},\mathbf{0}].$ (36)
It follows from (32) that the dynamic output feedback controller (35) is equal to the static state feedback controller $u(k)=-Kx(k)$, for all $k=0,1,2,\dots$. Thus, the proof is completed.
Combining Theorem 1 with the well-known linear quadratic regulator (4), the optimal dynamic output feedback controller can be given as
$u(k)$ $=-\mathcal{K}^{*}\eta(k)$ (37a)
$\eta(k+1)$ $=\mathcal{G}_{1}\eta(k)+\mathcal{G}_{2}\zeta(k)$ (37b)
with $\mathcal{K}^{*}=K^{*}M$.
3.2 Output Feedback Learning Control Approach
In this subsection, by only accessing the data information of input and output, we propose ADP-based learning algorithms to approximate the optimal feedback control gain $\mathcal{K}^{*}$, such that the optimality solution to the output feedback LQR problem, i.e., Problem 1, can be obtained. By combining the learning equation with the value iteration (VI) scheme and the policy iteration (PI) scheme, two model-free output feedback learning control approaches are proposed without any prior knowledge of system matrices.
To begin with, the well-known Bellman equation of the linear discrete-time systems can be written as
$V_{j}(x(k))=V_{j}(x(k+1))+\mathcal{C}_{x,u}(x(k),u(k))$ (38)
where $V_{j}(x(k))=x^{\mbox{\tiny T}}(k)P_{j}x(k)$ is the value function under the control policy $u(k)=-K_{j}x(k)$ and the utility function $\mathcal{C}_{x,u}(x(k),u(k))=x^{\mbox{\tiny T}}(k)Q_{x}x(k)+u^{\mbox{\tiny T}}(k)Ru(k)$ with the weight matrix $Q_{x}=Q_{x}^{\mbox{\tiny T}}\geq 0\in\mathbb{R}^{n\times n}$ and $R=R^{\mbox{\tiny T}}>0\in\mathbb{R}^{m\times m}$.
The model-based VI scheme and model-based PI scheme for estimating the optimal feedback control gain $K^{*}$ are recalled in Algorithms 1 and 2, respectively. Note that the utility function $\mathcal{C}_{y}(y(k),u(k))$ in (2) is a special case of $\mathcal{C}_{x,u}(x(k),u(k))$ by setting $Q_{x}=C^{\mbox{\tiny T}}Q_{y}C$.
Based on the Bellman equation (38), we next give an equivalent Bellman equation with the available state $\eta(k)$ of the internal model.
Along the system trajectories evaluated by (1), one can obtain
$x^{\mbox{\tiny T}}(k)P_{j}x(k)$ (39a)
$=$ $x^{\mbox{\tiny T}}(k\!+\!1)P_{j}x(k\!+\!1)+\mathcal{C}_{x,u}(x(k),u(k))$ (39b)
$=$ $\!\left[\!\!\begin{array}[]{cc}x(k)\\ u(k)\end{array}\!\!\right]^{\mbox{\tiny T}}\!\left[\!\!\begin{array}[]{cc}A^{\mbox{\tiny T}}P_{j}A\!+\!Q_{x}&A^{\mbox{\tiny T}}P_{j}B\\ B^{\mbox{\tiny T}}P_{j}A&B^{\mbox{\tiny T}}P_{j}B\!+\!R\end{array}\!\!\right]\!\left[\!\!\begin{array}[]{cc}x(k)\\ u(k)\end{array}\!\!\right].$ (39i)
For convenience of description, with the operators in Notation, define the functions $\delta_{v,w}(k)$ and $\delta_{v}(k)$ associated with the column vectors $v(k)$ and $w(k)$ as
$\delta_{v,w}(k)=$ $\left(w(k)\otimes v(k)\right)^{\mbox{\tiny T}}$ (40a)
$\delta_{v}(k)=$ $\mbox{vech}(2{v}(k){v}{{}^{\mbox{\tiny T}}}(k)-\mbox{dia}({v}(k))^{2})^{\mbox{\tiny T}}.$ (40b)
Define the data regressor functions $\Delta_{v}(k_{N})$, $\bar{\Delta}_{v}(k_{N})$ and $\Gamma_{v,w}(k_{N})$ associated with the column vectors sequences $v(k)$ and $w(k)$, $k=0,1,\dots,k_{N}$, as
${\Delta_{v}(k_{N})}=$ $[\delta_{v}(0)^{\mbox{\tiny T}},\delta_{v}(1)^{\mbox{\tiny T}},\dots,\delta_{v}(k_{N})^{\mbox{\tiny T}}]$ (41a)
$\bar{\Delta}_{v}(k_{N})=$ $[\bar{\delta}_{v}(0)^{\mbox{\tiny T}},\bar{\delta}_{v}(1)^{\mbox{\tiny T}},\dots,\bar{\delta}_{v}(k_{N})^{\mbox{\tiny T}}]$ (41b)
$\Gamma_{v,w}(k_{N})=$ $[\delta_{v,w}(0)^{\mbox{\tiny T}},\delta_{v,w}(1)^{\mbox{\tiny T}},\dots,\delta_{v,w}(k_{N})^{\mbox{\tiny T}}]$ (41c)
with $\bar{\delta}_{v}(k)=\delta_{v}(k+1)-\delta_{v}(k)$.
Algorithm 1. Model-Based VI Scheme
Initialization. Choose an initial positive semi-definite matrix $P_{0}$ and a small threshold $\epsilon>0$. Set $j=0$.
loop:
  Update $P_{j+1}$ from $P_{j+1}=A^{\mbox{\tiny T}}P_{j}A+Q_{x}-A^{\mbox{\tiny T}}P_{j}B\left(R+B^{\mbox{\tiny T}}P_{j}B\right)^{-1}B^{\mbox{\tiny T}}P_{j}A$ (42)   if $\|{P}_{j+1}-P_{j}\|<\epsilon$
    return $\hat{K}=\left(R+B^{\mbox{\tiny T}}P_{j}B\right)^{-1}B^{\mbox{\tiny T}}P_{j}A$
    break
  end if
  $j=j+1$
end loop
Algorithm 2. Model-Based PI Scheme
Initialization. Choose an initial admissible control gain $K_{0}$ and a small threshold $\epsilon>0$. Set $j=0$.
loop:
  Solve $P_{j}$ from $\mathbf{0}=(A-BK_{j})^{\mbox{\tiny T}}P_{j}(A-BK_{j})-P_{j}+Q_{x}+K_{j}^{\mbox{\tiny T}}RK_{j}$ (43)   if $j\geq 1$ and $\|{P}_{j}-P_{j-1}\|<\epsilon$
    return $\hat{K}=K_{j}$
    break
  else
    Update $K_{j+1}$ from $K_{j+1}=\left(R+B^{\mbox{\tiny T}}P_{j}B\right)^{-1}B^{\mbox{\tiny T}}P_{j}A$ (44)   end if
  $j=j+1$
end loop
By the equivalence relationship given in (32), substituting the collected data vectors $\eta(k)$ and $u(k)$ into the functions $\delta_{v,w}$ and $\delta_{v}$, (39) can be rewritten as
$\eta^{\mbox{\tiny T}}(k)\mathcal{P}_{j}\eta(k)$ (45a)
$=$ $\eta^{\mbox{\tiny T}}(k+1)\mathcal{P}_{j}\eta(k+1)+\mathcal{C}_{\eta,u}(\eta(k),u(k))$ (45b)
$=$ $\varphi(k)\left[\begin{matrix}\mbox{vech}(\mathcal{P}_{\eta,\eta}^{(j)})\\ \mbox{vec}({\mathcal{P}}_{\eta,u}^{(j)})\\ \mbox{vech}({\mathcal{P}}_{u,u}^{(j)})\end{matrix}\right]$ (45c)
where
${\mathcal{P}}_{j}=$ $M^{\mbox{\tiny T}}P_{j}M$
$\mathcal{P}_{\eta,\eta}^{(j)}=$ $M^{\mbox{\tiny T}}(A^{\mbox{\tiny T}}P_{j}A+Q_{x})M$
$\mathcal{P}_{\eta,u}^{(j)}=$ $M^{\mbox{\tiny T}}A^{\mbox{\tiny T}}P_{j}B$
$\mathcal{P}_{u,u}^{(j)}=$ $(B^{\mbox{\tiny T}}P_{j}B+R)$
and the data regressors are
$\varphi(k)=$ $\left[{\delta_{\eta}}(k),~{}{2{\delta_{{\eta},{Ru}}}(k)},~{}{\delta_{u}}(k)\right]$
$\mathcal{C}_{\eta,u}(\eta(k),u(k))=$ $\eta^{\mbox{\tiny T}}(k)M^{\mbox{\tiny T}}Q_{x}M\eta(k)+u^{\mbox{\tiny T}}(k)Ru(k).$
Under the utility function defined in (2) which is the case of $Q_{x}=C^{\mbox{\tiny T}}Q_{y}C$, the data regressor $\mathcal{C}_{\eta,u}(\eta(k),u(k))$ is calculable and given by
$\mathcal{C}_{\eta,u}(\eta(k),u(k))\triangleq$ $\mathcal{C}_{y}(y(k),u(k))$
$=$ $y^{\mbox{\tiny T}}(k)Q_{y}y(k)+u^{\mbox{\tiny T}}(k)Ru(k).$ (46)
Based on the data-based Bellman equation (45), the model-based VI scheme (Algorithm 1), and the model-based PI scheme (Algorithm 2), we next establish the model-free VI scheme and model-free PI scheme for solving Problem 1.
3.2.1 Model-Free VI Scheme
Given an initial symmetric positive definite matrix $\mathcal{P}_{0}\in\mathbb{R}^{n_{z}\times n_{z}}$, for each iteration $j=0,1,2,\dots$, the equality between (45b) and (45c) always holds for any utility function $\mathcal{C}_{\eta,u}(\eta(k),u(k))$. Thus, the vector-formed data-based learning equation is given as
$\varphi(k)\left[\begin{matrix}\mbox{vech}(\mathcal{P}_{\eta,\eta}^{(j)})\\ \mbox{vec}({\mathcal{P}}_{\eta,u}^{(j)})\\ \mbox{vech}({\mathcal{P}}_{u,u}^{(j)})\end{matrix}\right]=\psi(k)\mbox{vech}\left({\mathcal{P}}_{j}\right)+\mathcal{C}(k)$ (47)
where $\psi(k)\!=\!{\delta_{\eta}}(k+1)$ and $\mathcal{C}(k)=\mathcal{C}_{\eta,u}(\eta(k),u(k))$ is the value of utility function at instant $k$.
By the least squares implementation, the matrix-formed data-based learning equation used to estimate the unknown parameters $\left(\mathcal{P}_{\eta,\eta}^{(j)},~{}\mathcal{P}_{\eta,u}^{(j)},~{}\mathcal{P}_{u,u}^{(j)}\right)$ is given as
$\Pi(k_{N})\left[\begin{matrix}\mbox{vech}(\mathcal{P}_{\eta,\eta}^{(j)})\\ \mbox{vec}({\mathcal{P}}_{\eta,u}^{(j)})\\ \mbox{vech}({\mathcal{P}}_{u,u}^{(j)})\end{matrix}\right]=\Psi(k_{N})\mbox{vech}\left({\mathcal{P}}_{j}\right)+\mathcal{L}(k_{N})$ (48)
where $\Pi(k_{N})={\mbox{col}}\left(\varphi(0),\varphi(1),\dots,\varphi(k_{N})\right)\in\mathbb{R}^{N\times n_{\varphi}}$ and $\Psi(k_{N})={\mbox{col}}\left(\psi(0),\psi(1),\dots,\psi(k_{N})\right)\in\mathbb{R}^{N\times n_{\psi}}$ with dimensions $n_{\varphi}=\frac{(n_{z}+m)(n_{z}+m+1)}{2}$ and $n_{\psi}=\frac{n_{z}(n_{z}+1)}{2}$, and $\mathcal{L}(k_{N})={\mbox{col}}\left(\mathcal{C}\left(0\right),\mathcal{C}\left(1\right),\dots,\mathcal{C}\left(k_{N}\right)\right)$. The matrices $\Pi(k_{N})$, $\Psi(k_{N})$, and $\mathcal{L}(k_{N})$ are the data regressors during the data collection process $k\in[k_{0},k_{N}]$.
Combining the model-based VI scheme with the solution to (48), the value evaluation equation is given by
$\mathcal{P}_{j+1}\!=\!\mathcal{P}_{\eta,\eta}^{(j)}\!-\!\mathcal{P}_{\eta,u}^{(j)}\left(\mathcal{P}_{u,u}^{(j)}\right)^{-1}\left(\mathcal{P}_{\eta,u}^{(j)}\right)^{\mbox{\tiny T}}.$ (49)
The control policy can be estimated by
$\mathcal{K}_{j}=\left(\mathcal{P}_{u,u}^{(j)}\right)^{-1}\left(\mathcal{P}_{\eta,u}^{(j)}\right)^{\mbox{\tiny T}}.$ (50)
Before we present the model-free learning algorithms with the VI scheme, we give a rank condition that is used to determine whether the collected data is enough for implementing the learning process.
Theorem 2
If there exists an instant $k_{N}^{*}$ such that the rank condition
$\textnormal{rank}\left(\Pi(k_{N}^{*})\right)=\frac{(n_{z}+m)(n_{z}+m+1)}{2}$ (51)
is satisfied, then, for every iteration $j=0,1,2,\dots$, the VI-based learning equation (48) has a unique solution.
Proof. If the rank condition (51) is satisfied at $k_{N^{*}}$, then $\Pi(k_{N}^{*})$ in (48) is of full column rank at each iteration $j$, since $\Pi(k_{N}^{*})$ only relies on the collected data and is independent on $\mathcal{P}_{j}$ and $\mathcal{K}_{j}$. Thus, there always exists a unique solution to the VI-based learning equation (48) for every iteration $j$.
Finally, the proposed output feedback learning control approach with the VI scheme can be summarized in Algorithm 3.
Algorithm 3. Model-Free VI-Based Learning Control
Initialization. Choose a symmetric positive semi-definite matrix $\mathcal{P}_{0}$. Select an pair $(\mathcal{G}_{1},\mathcal{G}_{2})$, an initial internal state $\eta(0)$, an initial control gain ${\mathcal{K}_{0}}$, and a small threshold $\epsilon>0$. Apply the initial control policy $u_{0}=-{\mathcal{K}_{0}}{\eta}+\xi$ with exploration noise $\xi$. Set $j=0$.
Data Collection. Collect the sampled data $(u(k),y(k))$ and calculate $(\eta(k),\eta(k+1))$ until the rank condition (51) holds.
loop:
  Calculate $\mathcal{P}_{\eta,\eta}^{(j)}$, $\mathcal{P}_{\eta,u}^{(j)}$, and $\mathcal{P}_{u,u}^{(j)}$ by (48)
  Update $\mathcal{P}_{j+1}$ by (49)
  if $\|\mathcal{P}_{j+1}-\mathcal{P}_{j}\|<\epsilon$
    Calculate $\mathcal{K}_{j}$ by (50)
    return $\tilde{\mathcal{K}}^{*}=\mathcal{K}_{j}$
    break
  end if
  $j=j+1$
end loop
3.2.2 Model-Free PI Scheme
Given an initial admissible control gain $\mathcal{K}_{0}$, for every iteration $j=0,1,2,\dots$, one can obtain
$~{}\eta^{\mbox{\tiny T}}(k+1)\mathcal{P}_{j}\eta(k+1)-\eta^{\mbox{\tiny T}}(k)\mathcal{P}_{j}\eta(k)$
$=$ $~{}2\eta^{\mbox{\tiny T}}(k)M^{\mbox{\tiny T}}(A-BK_{j})^{\mbox{\tiny T}}P_{j}B\left(u(k)+\mathcal{K}_{j}\eta(k)\right)$
$+\left(u(k)+\mathcal{K}_{j}\eta(k)\right)^{\mbox{\tiny T}}B^{\mbox{\tiny T}}P_{j}B\left(u(k)+\mathcal{K}_{j}\eta(k)\right)$
$-y^{\mbox{\tiny T}}(k)Q_{y}y(k)-\eta^{\mbox{\tiny T}}(k)\mathcal{K}_{j}^{\mbox{\tiny T}}R\mathcal{K}_{j}\eta(k).$ (52)
Define $\hat{u}_{j}(k)=u(k)+\mathcal{K}_{j}\eta(k)$. Substituting the data $\eta(k)$ and $\hat{u}_{j}(k)$ into the functions defined in (40), according to (3.2.2), the vector-formed PI-based learning equation is given as
${\hat{\varphi}}_{j}(k)\left[\begin{matrix}\mbox{vech}(\mathcal{P}_{j})\\ \mbox{vec}({\mathcal{P}}_{A,B}^{(j)})\\ \mbox{vech}({\mathcal{P}}_{B,B}^{(j)})\end{matrix}\right]=\phi(k)\left[\begin{matrix}\mbox{vech}(Q_{y})\\ \mbox{vech}(\left(\mathcal{K}_{j}^{\mbox{\tiny T}}R\mathcal{K}_{j}\right)\end{matrix}\right]$ (53)
with
${\hat{\varphi}}_{j}(k)=$ $\left[-\bar{\delta}_{\eta}(k),~{}{2{\delta_{{\eta},{R\hat{u}_{j}}}}(k)},~{}{\delta_{\hat{u}_{j}}}(k)\right]$
$\phi(k)=$ $\left[{\delta_{y}}(k),~{}{\delta_{\eta}(k)}\right]$
$\mathcal{P}_{A,B}^{(j)}=$ $M^{\mbox{\tiny T}}(A-BK_{j})^{\mbox{\tiny T}}P_{j}B$
$\mathcal{P}_{B,B}^{(j)}=$ $B^{\mbox{\tiny T}}P_{j}B.$
By using the least squares method, the matrix-formed data-based learning equation with the PI scheme is given as
${\hat{\Pi}}_{j}(k_{N})\left[\begin{matrix}\mbox{vech}(\mathcal{P}_{j})\\ \mbox{vec}({\mathcal{P}}_{A,B}^{(j)})\\ \mbox{vech}({\mathcal{P}}_{B,B}^{(j)})\end{matrix}\right]=\Phi(k_{N})\left[\begin{matrix}\mbox{vech}(Q_{y})\\ \mbox{vech}(\left(\mathcal{K}_{j}^{\mbox{\tiny T}}R\mathcal{K}_{j}\right)\end{matrix}\right]$ (54)
where ${\hat{\Pi}}_{j}(k_{N})={\mbox{col}}\left({\hat{\varphi}}_{j}(0),{\hat{\varphi}}_{j}(1),\dots,{\hat{\varphi}}_{j}(k_{N})\right)\in\mathbb{R}^{N\times n_{\varphi}}$ and $\Phi(k_{N})={\mbox{col}}\left(\phi(0),\phi(1),\dots,\phi(k_{N})\right)\in\mathbb{R}^{N\times n_{\phi}}$ with dimension $n_{\phi}=\frac{(n_{z}+p)(n_{z}+p+1)}{2}$.
Combining the model-based PI scheme with the solution to (54), $\mathcal{P}_{j}$ can be estimated by calculating (54), and the control policy can be updated by
$\mathcal{K}_{j+1}=\left(R+\mathcal{P}_{B,B}^{(j)}\right)^{-1}\left(\left(\mathcal{P}_{A,B}^{(j)}\right)^{\mbox{\tiny T}}+\mathcal{P}_{B,B}^{(j)}\mathcal{K}_{j}\right).$ (55)
Algorithm 4. Model-Free PI-Based Learning Control
Initialization. Choose an initial admissible control gain $\mathcal{K}_{0}$. Select an pair $(\mathcal{G}_{1},\mathcal{G}_{2})$, an initial internal state $\eta(0)$, and a small threshold $\epsilon>0$. Apply the initial control policy $u_{0}=-{\mathcal{K}_{0}}{\eta}+\xi$ with exploration noise $\xi$. Set $j=0$.
Data Collection. Collect the sampled data $(u(k),y(k))$ and calculate $(\eta(k),\eta(k+1))$ until the rank condition (51) holds.
loop:
  Calculate $\mathcal{P}_{j}$, $\mathcal{P}_{A,B}^{(j)}$, and $\mathcal{P}_{B,B}^{(j)}$ by (54).
  if $j\geq 1$ and $\|\mathcal{P}_{j}-\mathcal{P}_{j-1}\|<\epsilon$
    return $\tilde{\mathcal{K}}^{*}=\mathcal{K}_{j}$
    break
  else
    Update $\mathcal{K}_{j+1}$ by (55).
  end if
  $j=j+1$
end loop
We next show that the same rank condition (51) can be used to determine whether the collected data is enough for implementing the learning process of the model-free PI scheme.
Theorem 3
If there exists an instant $k_{N}^{*}$ such that the rank condition (51) is satisfied, then, for every iteration $j=0,1,2,\dots$, the PI-based learning equation (54) has a unique solution.
Proof. We give a proof by contradiction to show the existence of a unique solution to (54) for every iteration. Suppose that ${\hat{\Pi}}_{j}(k_{N}^{*})$ in (54) is not of full column rank such that (51) is satisfied. This implies that
${\hat{\Pi}}_{j}(k_{N}^{*})\left[\begin{matrix}\mbox{vech}(\mathcal{P}_{j})\\ \mbox{vec}({\mathcal{P}}_{A,B}^{(j)})\\ \mbox{vech}({\mathcal{P}}_{B,B}^{(j)})\end{matrix}\right]=\mathbf{0}$ (56)
has a nonzero solution denoted by $\mathbb{W}={\mbox{col}}\left(V,~{}W_{1},~{}W_{2}\right)$ with $V=\mbox{vech}({M^{\mbox{\tiny T}}V^{m}M})$, $W_{1}=\mbox{vec}({M^{\mbox{\tiny T}}W_{1}^{m}})$ and $W_{2}=\mbox{vech}({W_{2}^{m}})$. Substituting the data $\eta(k)$, $u(k)$ and $\hat{u}_{j}(k)$ into the data regressor functions defined in (41), it follows from (56) that
${\bar{\Delta}_{\eta}}V-2\Gamma_{\eta,\hat{u}_{j}}W_{1}-\Delta_{\hat{u}_{j}}W_{2}=\mathbf{0}.$ (57)
By (1a) and (32), the first part in the left-hand side of (57) becomes
${\bar{\Delta}_{\eta}}V\!=$ $\Delta_{\eta}\mbox{vech}\left(M^{\mbox{\tiny T}}\left(\left(A\!-\!BK_{j}\right)^{\mbox{\tiny T}}\!V^{m}\!\left(A\!-\!BK_{j}\right)\!-\!V^{m}\right)M\right)$
$+\Delta_{\eta}\mbox{vech}\left(M^{\mbox{\tiny T}}\left(2A^{\mbox{\tiny T}}V^{m}BK_{j}\!-\!K_{j}^{\mbox{\tiny T}}B^{\mbox{\tiny T}}V^{m}BK_{j}\right)M\right)$
$+2\Gamma_{\eta,u}\mbox{vec}\left(M^{\mbox{\tiny T}}(A^{\mbox{\tiny T}}V^{m}B)\right)\!+\!\Delta_{u}\mbox{vech}(B^{\mbox{\tiny T}}V^{m}B).$ (58)
Combining (57) and (58), we have
${\Delta_{{\eta}}}\mbox{vech}(M^{\mbox{\tiny T}}{\Omega_{1}}M)\!+\!2{\Gamma_{{\eta},{u}}}\mbox{vec}(M^{\mbox{\tiny T}}{\Omega_{2}})\!+\!{\Delta_{{u}}}\mbox{vech}({\Omega_{3}})\!=\!\mathbf{0}$ (59)
where
${\Omega_{1}}=$ $\left(A\!-\!BK_{j}\right)^{\mbox{\tiny T}}\!V^{m}\!\left(A\!-\!BK_{j}\right)-V^{m}$
$+2\left(A^{\mbox{\tiny T}}V^{m}B-W_{1}^{m}-\left(K_{j}\right)^{\mbox{\tiny T}}W_{2}^{m}\right)K_{j}$ (60a)
$-\left(K_{j}\right)^{\mbox{\tiny T}}\left(B^{\mbox{\tiny T}}V^{m}B-W_{2}^{m}\right)K_{j}$
${\Omega_{2}}=$ $A^{\mbox{\tiny T}}V^{m}B-W_{1}^{m}-\left(K_{j}\right)^{\mbox{\tiny T}}W_{3}^{m}$ (60b)
${\Omega_{3}}=$ $B^{\mbox{\tiny T}}V^{m}B-W_{2}^{m}.$ (60c)
If the rank condition (51) is satisfied, then (59) only has zero solution, i.e., ${\Omega_{1}}=\mathbf{0}$, ${\Omega_{2}}=\mathbf{0}$ and ${\Omega_{3}}=\mathbf{0}$. Then, we have
$\begin{array}[]{l}\left(A\!-\!BK_{j}\right)^{\mbox{\tiny T}}\!V^{m}\!\left(A\!-\!BK_{j}\right)-V^{m}=\mathbf{0}.\end{array}$ (62)
According to Hewer (1971) that the matrix $\left(A-BK_{j}\right)$ is Schur for all $j=0,1,2,\dots$, the only solution to (62) is ${V^{m}}=\mathbf{0}$. Substituting $V^{m}=\mathbf{0}$ into (60), we have ${W_{1}^{m}}=\mathbf{0}$, ${W_{2}^{m}}=\mathbf{0}$, and ${W_{3}^{m}}=\mathbf{0}$, which contradicts the fact that $\mathbb{W}$ is a nonzero solution to (56). Thus, if (51) holds, then, for all $j=0,1,2,\dots$, ${\hat{\Pi}}_{j}(k_{N}^{*})$ is of full rank.
Finally, the proposed output feedback learning control approach with the PI scheme is summarized in Algorithm 4.
4 Convergence, Stability, and Optimality Analysis
In this section, we give a detailed analysis of convergence, stability, and optimality of the proposed output feedback learning control approaches and provide a model-free stability criterion. Moreover, we propose a switched iteration scheme integrated with PI and VI by leveraging the stability criterion, which can relax the requirement of an initial admissible controller in the PI scheme and accelerate the convergence rate in the VI scheme.
4.1 Convergence analysis
To show the convergence of the learning algorithm, we first give a rank condition for the parameterization matrix $M$ as follows:
Assumption 2
The parameterization matrix $M$ satisfies
$\textnormal{rank}(M)=n.$ (63)
Remark 1
In Rizvi & Lin (2023), a necessary condition for the convergence of learning algorithms with both PI and VI is given. That is, the parameterization matrix $\bar{W}$ in (11) is of full rank. It is noted that the state observer error parameterization matrix $M_{\varepsilon_{x}\leftarrow\eta_{\varepsilon}}$ is constructed in $M$. Thus, compared with the full rank condition of $\bar{W}$, condition (63) is easier to hold.
We next show the full rank (63) in Lemma 1 can be guaranteed by providing a simple design method.
Lemma 1
If the matrix $A_{L}$ has $n$ distinct eigenvalues, then there exists a square matrix $M_{\varepsilon_{x}\leftarrow\eta_{\varepsilon}}$ that is of full rank.
Proof. It follows from (19) that the square matrix $M_{\varepsilon_{x}\leftarrow\eta_{\varepsilon}}=M_{\varepsilon_{x}}M_{\eta_{\varepsilon}}^{-1}$ is of full rank if and only if $M_{\varepsilon_{x}}$ is non-singular, since $M_{\eta_{\varepsilon}}$ is a user-defined non-singular matrix. Let $\varepsilon_{0}=\varepsilon_{x}(0)$ be the initial observer error. Then, we give the existence proof of $M_{\varepsilon_{x}}$ being of full rank as follows.
The determinant and adjoint matrix of $(zI_{n}-A_{L})$ can be rewritten as
$\mbox{det}(zI_{n}\!-\!A_{L})$ $\!=\!z^{n}\!+\!b_{1}z^{n\!-\!1}\!+\!b_{2}z^{n\!-\!2}\!+\!\cdots\!+\!b_{n-1}z\!+\!b_{n}$ (64)
$\mbox{adj}(zI_{n}\!-\!A_{L})$ $\!=\!\Upsilon_{1}z^{n\!-\!1}+\Upsilon_{2}z^{n\!-\!2}+\cdots+\Upsilon_{n\!-\!1}z\!+\!\Upsilon_{n}$ (65)
where $b_{i}\in\mathbb{R}$ and $\Upsilon_{i}\in\mathbb{R}^{n}$, $i=1,2,\dots,{n}$, are the constant coefficients and matrix coefficients, respectively. According to the properties of the determinant and adjoint matrix, one can obtain
$\mbox{det}(zI_{n}-A_{L})I_{n}$
$=$ $\mbox{adj}(zI_{n}-A_{L})(zI_{n}-A_{L})$
$=$ $\Upsilon_{1}z^{n}+(\Upsilon_{2}-\Upsilon_{1}A_{L})z^{n-1}+\cdots+(\Upsilon_{i+1}-\Upsilon_{i}A_{L})z^{n-i}$
$+\cdots+(\Upsilon_{n}-\Upsilon_{n-1}A_{L})z-\Upsilon_{n}A_{L}.$ (66)
From the equality of the coefficients of (64)–(4.1), we have
$\Upsilon_{i+1}=\Upsilon_{i}A_{L}+b_{i}I_{n},~{}~{}~{}~{}i=1,2,\dots,n-1$
with $\Upsilon_{1}=I_{n}$.
Then, $M_{\varepsilon_{x}}$ can be represented by
$M_{\varepsilon_{x}}=$ $\left[\Upsilon_{n}\varepsilon_{0},~{}\Upsilon_{n-1}\varepsilon_{0},~{}\dots~{},~{}\Upsilon_{3}\varepsilon_{0},~{}\Upsilon_{2}\varepsilon_{0},~{}\Upsilon_{1}\varepsilon_{0},\right]$
$=$ $\left[A_{L}^{n-1}\varepsilon_{0}+b_{1}A_{L}^{n-2}\varepsilon_{0}+\cdots+b_{n-1}\varepsilon_{0},~{}\dots,\right.$
$~{}\left.~{}\dots,A_{L}^{2}\varepsilon_{0}+b_{1}A_{L}\varepsilon_{0}+b_{2}\varepsilon_{0},~{}A_{L}\varepsilon_{0}+b_{1}\varepsilon_{0},~{}\varepsilon_{0}\right].$
Thus, we have
$\mbox{rank}\left(M_{\varepsilon_{x}}\right)=\mbox{rank}\left(\left[\varepsilon_{0},~{}A_{L}\varepsilon_{0},~{}A_{L}^{2}\varepsilon_{0},~{}\dots,~{}A_{L}^{n-1}\varepsilon_{0}\right]\right).$
If the matrix $A_{L}$ has $n$ distinct eigenvalues, i.e., $\lambda_{1},\lambda_{2},\dots,\lambda_{n}$ are not equal to each other, then there exists a non-singular matrix $U$ which can transform $A_{L}$ into a diagram matrix $\mathcal{J}_{L}=\mbox{block diag}[\lambda_{1},\lambda_{2},\dots,\lambda_{n}]$, i.e., $U^{-1}A_{L}U=\mathcal{J}_{L}$. Letting $\bar{\varepsilon}_{0}=U^{-1}\varepsilon_{0}$, we have
$U^{-1}$ $\left[\varepsilon_{0},~{}A_{L}\varepsilon_{0},~{}A_{L}^{2}\varepsilon_{0},~{}\cdots,~{}A_{L}^{n-1}\varepsilon_{0}\right]$
$=$ $\left[\bar{\varepsilon}_{0},~{}U^{-1}A_{L}U\bar{\varepsilon}_{0},~{}U^{-1}A_{L}^{2}U\bar{\varepsilon}_{0},~{}\cdots,~{}U^{-1}A_{L}^{n-1}U\bar{\varepsilon}_{0}\right]$
$=$ $\left[\bar{\varepsilon}_{0},~{}\mathcal{J}_{L}\bar{\varepsilon}_{0},~{}\mathcal{J}_{L}^{2}\bar{\varepsilon}_{0},~{}\cdots,~{}\mathcal{J}_{L}^{n-1}\bar{\varepsilon}_{0}\right]$
$=$ $\mbox{dia}(\bar{\varepsilon}_{0})\left[\begin{matrix}1&\lambda_{1}&\lambda_{1}^{2}&\cdots&\lambda_{1}^{n-1}\\ 1&\lambda_{2}&\lambda_{2}^{2}&\cdots&\lambda_{2}^{n-1}\\ \vdots&\vdots&\vdots&\cdots&\vdots\\ 1&\lambda_{n}&\lambda_{n}^{2}&\cdots&\lambda_{n}^{n-1}\\ \end{matrix}\right].$
If all the elements of $\bar{\varepsilon}_{0}$ are nonzero, the matrix $M_{\varepsilon_{x}}$ is of full rank. Therefore, since $M_{\eta_{\varepsilon}}$ is chosen as non-singular, there exists a matrix $M_{\varepsilon_{x}\leftarrow\eta_{\varepsilon}}=M_{\varepsilon_{x}}M_{\eta_{\varepsilon}}^{-1}$ that is of full rank. This completes the proof.
Remark 2
In some existing output feedback learning algorithms, for instance, Rizvi & Lin (2023, 2020); Chen et al. (2023); Gao et al. (2016), the constructed parameterization matrices are in the form of either $\bar{M}$ in (2.2) or $\bar{W}$ in (11). Under Assumption 1, it follows from (2.2) that the matrix $\bar{M}$ may not be of full row rank when the system matrix $A$ has the eigenvalue 00, and the matrix $\bar{W}$ may not be of full row rank when the observer matrices $(A-LC)$ and $A$ have common eigenvalues according to Rizvi & Lin (2023, Theorem 4). Different from the parameterization matrices $\bar{M}$ and $\bar{W}$, the rank condition of parameterization matrix $M$ in this paper can be satisfied by choosing the constructed $M_{\varepsilon_{x}\leftarrow z_{\varepsilon}}$ to be full rank. According to Lemma 1, there exists matrix $M_{\varepsilon_{x}\leftarrow z_{\varepsilon}}$ that has full rank by setting all the eigenvalues of $A_{L}$ distinct. This pole assignment can be easily achieved which only requires Assumption 1 satisfied.
Remark 3
The exploration noise is usually added to the control input during the data collection process to ensure the rank condition (51), for instance, Bian & Jiang (2016); Jiang & Jiang (2012); Modares et al. (2016); Rizvi & Lin (2023). Thus, it is not difficult to get a nonzero state $U^{-1}{\varepsilon}_{x}(k)$ at an instant $k$ that can be set as the initial instant 00 for the learning process. Furthermore, we provide a way to check if the rank condition (63) of $M$ is met or not in Subsection 4.2.
Now, we are ready to give the convergence analysis of Algorithms 3 and 4 shown in the following Theorems 4 and 5, respectively.
Theorem 4
For any given initial positive semi-definite symmetric matrix $\mathcal{P}_{0}$, if the rank conditions (51) and (63) hold, the control gain $\mathcal{K}_{j}$ estimated by the model-free VI-based Algorithm 3 converges to the optimal value $\mathcal{K}^{*}$, i.e., $\lim_{j\rightarrow\infty}\mathcal{K}_{j}=\mathcal{K}^{*}$.
Proof. By the solution to the iterative learning equation (48), the value evaluation equation (49) can be written as
$\mathcal{P}_{j+1}=M^{\mbox{\tiny T}}\left(A^{\mbox{\tiny T}}P_{j}A+C^{\mbox{\tiny T}}Q_{y}C-A^{\mbox{\tiny T}}P_{j}B\left(R+B^{\mbox{\tiny T}}P_{j}B\right)^{-1}B^{\mbox{\tiny T}}P_{j}A\right)M$
If the rank condition (63) holds, (49) is equivalent to (42). As proven in Dorato & Levis (1971), for any given $P_{0}=P_{0}^{\mbox{\tiny T}}\geq 0$, the $P_{j}$ calculated by (42) converges to $\bar{P}^{*}$ as $j\rightarrow\infty$. In Algorithm 3, the value evaluation equation (49) with the solution to the iterative learning equation (48) is transformed from (42). As shown in Theorem 2, the existence of the unique solution of (48) can be ensured once the rank condition (51) is satisfied. Thus, in Algorithm 3, for any given ${\mathcal{P}_{0}=\mathcal{P}_{0}^{\mbox{\tiny T}}\geq 0}$, substituting ${\mathcal{P}_{j}}$ into (48) for each iteration $j=0,1,\dots$, ${\mathcal{P}_{j+1}}$ can be recalculated uniquely from (49). Therefore, as $j\rightarrow\infty$, $\mathcal{P}_{j+1}$ converges to ${\mathcal{P}^{*}}$, which implies that $\mathcal{K}_{j}$ converges to $\mathcal{K}^{*}$. Thus, the proof is completed.
Theorem 5
For any given initial admissible control policy $\mathcal{K}_{0}$, if the rank conditions (51) and (63) hold, the control gain $\mathcal{K}_{j}$ estimated by the model-free PI-based Algorithm 4 converges to the optimal value $\mathcal{K}^{*}$, i.e., $\lim_{j\rightarrow\infty}\mathcal{K}_{j}=\mathcal{K}^{*}$.
Proof. Along the similar way as the proof in Theorem 4, when the rank conditions (51) and (63) are satisfied, it is not hard to have $\lim_{j\rightarrow\infty}\mathcal{K}_{j}=\mathcal{K}^{*}$ with Algorithm 4 by the convergence of the PI scheme proven in Hewer (1971). Thus, the proof is omitted.
Remark 4
Different from the parameterization matrix constructed in Rizvi & Lin (2023); Chen et al. (2022), we develop a method to construct a square matrix $M_{\varepsilon_{x}\leftarrow\eta_{\varepsilon}}$ that is of full rank. It can not only increase the flexibility of the matrix $M$ but also enable us to analyze the stability of the closed-loop system with the estimated control policy at any iteration $j$ shown in Subsection 4.2.
4.2 Stability Analysis
By using a stopping criterion with a small threshold $\epsilon$, we can obtain an estimated control gain $\tilde{\mathcal{K}}^{*}={\tilde{K}}^{*}M$ by Algorithm 3 or Algorithm 4. For the PI-based scheme, as proven in Hewer (1971), the estimated control policy has the property that $\sigma(A-B{\tilde{K}}^{*})\in\mathbb{U}^{-}$ for $j=0,1,2,\dots$. This indicates that the stability of the closed-loop system with the PI-based scheme can be guaranteed at every iteration $j$.
However, for the VI-based scheme, this guaranteed stability property does not hold. Moreover, it is important to notice that the matrix $\left(C^{\mbox{\tiny T}}Q_{y}C+(K^{*})^{\mbox{\tiny T}}RK^{*}\right)$ is positive semi-definite in some cases, for example, the dimension of the system satisfies $n>p+m$. In these cases, the estimated matrix $\left(\left(A-B{\tilde{K}}^{*}\right)^{\mbox{\tiny T}}\hat{P}^{*}\left(A-B{\tilde{K}}^{*}\right)-\tilde{P}^{*}\right)$ by using the stopping criterion may not be positive semi-definite. This implies that the closed-loop system with the estimated control gain $\tilde{\mathcal{K}}^{*}$ by Algorithm 3 may not achieve globally asymptotically stable at the origin. Thus, it is worth providing a stability criterion for the closed-loop system with the proposed output feedback learning control approach.
To analyze the stability of the linear discrete-time closed-loop systems, we recall the classical Lyapunov theorem.
Lemma 2
Chen (1999) Given any positive definite symmetric matrix $Q\in\mathbb{R}^{n\times n}$, there exists a unique positive definite symmetric solution $P$ to the Lyapunov equation
$\left(A-BK\right)^{\mbox{\tiny T}}P\left(A-BK\right)-P=-Q$
if and only if $\sigma(A-BK)\in\mathbb{U}^{-}$.
In the case of completely unknown system matrices, an alternative way to analyze the stability of the closed-loop systems is to construct an equivalent Lyapunov equation by using the data information. Although the data of states, as the most important information that can reflect system dynamics, is unmeasurable, it is fortunate that there always exists a linear relationship between the designed internal state $\eta$ and the real state $x$ as shown in (32). By using the solution to an auxiliary data-based equation, we next provide a model-free stability criterion for the closed-loop system under the estimated control policy at any iteration $j$.
Define the auxiliary data-based equation as
$\varphi(k)\left[\begin{matrix}\mbox{vech}(\bar{\mathcal{P}}_{\eta,\eta}^{(j)})\\ \mbox{vec}({\mathcal{P}}_{\eta,u}^{(j)})\\ \mbox{vech}({\mathcal{P}}_{u,u}^{(j)})\end{matrix}\right]=\psi(k)\mbox{vech}\left({\mathcal{P}}_{j}\right)+\bar{\mathcal{C}}(k)$ (67)
with $\bar{\mathcal{P}}_{\eta,\eta}^{(j)}=\mathcal{P}_{\eta,\eta}^{(j)}=M^{\mbox{\tiny T}}A^{\mbox{\tiny T}}P_{j}AM$ and $\bar{\mathcal{C}}(k)=u^{\mbox{\tiny T}}(k)Ru(k)$, which is the modification of (47) with utility function $\bar{\mathcal{C}}(k)$. By the solution to (67), we define
${H}_{j}(P_{j},K_{j})=\bar{\mathcal{P}}_{\eta,\eta}^{(j)}-2{\mathcal{P}}_{\eta,u}^{(j)}\mathcal{K}_{j}\!+\!\left(\mathcal{K}_{j}\right)^{\mbox{\tiny T}}{\mathcal{P}}_{u,u}^{(j)}\mathcal{K}_{j}-\mathcal{P}_{j}$
where $\mathcal{K}_{j}=\left(\mathcal{P}_{u,u}^{(j)}\right)^{-1}\left(\mathcal{P}_{\eta,u}^{(j)}\right)^{\mbox{\tiny T}}$. Then, by extracting the last $n\times n$ submatrix from ${H}_{j}(P_{j},K_{j})$ to form a new matrix $\mathcal{H}_{j}(P_{j},K_{j})$, one can obtain
$\mathcal{H}_{j}(P_{j},K_{j})\!=M_{\varepsilon_{x}\leftarrow\eta_{\varepsilon}}^{\mbox{\tiny T}}\left(\left(A\!-\!BK_{j}\right)^{\mbox{\tiny T}}\!P_{j}\!\left(A\!-\!BK_{j}\right)\!-\!P_{j}\right)M_{\varepsilon_{x}\leftarrow\eta_{\varepsilon}}.$
Note that the non-singularity of $M_{\varepsilon_{x}\leftarrow\eta_{\varepsilon}}$ can be easily verified in the learning process by checking $\mbox{det}(\mathcal{H}_{j}(K_{j},P_{j}))\neq 0$.
Then, we provide a way to guarantee the stability of the closed-loop system with the estimated control policy by VI-based model-free algorithm.
Lemma 3
Given any symmetric positive definite matrix $Q_{\epsilon}$, under rank condition (51) and (63), then there always exists an iteration $j$ satisfying the stability criterion
$\mathcal{H}_{j}(P_{j},K_{j})<0$ (68)
by employing the model free VI-based scheme with the utility function
$\tilde{\mathcal{C}}(k)=\eta^{\mbox{\tiny T}}(k)Q_{\epsilon}\eta(k)+y^{\mbox{\tiny T}}(k)Q_{y}y(k)+u^{\mbox{\tiny T}}(k)Ru(k).$ (69)
Proof. We give a proof by contradiction for the existence of the iteration $j$. Let $\bar{Q}_{x}=(MM^{\mbox{\tiny T}})^{-1}MQ_{\varepsilon}M^{\mbox{\tiny T}}(MM^{\mbox{\tiny T}})^{-1}$ and ${\bar{Q}}_{\varepsilon}=Q_{\varepsilon}-M^{\mbox{\tiny T}}{\bar{Q}}_{x}M$. The utility function (69) can be rewritten as
$\tilde{\mathcal{C}}(k)=\eta^{\mbox{\tiny T}}(k){\bar{Q}}_{\epsilon}\eta(k)+x^{\mbox{\tiny T}}(k){\tilde{Q}}_{x}x(k)+u^{\mbox{\tiny T}}(k)Ru(k)$
with ${\tilde{Q}}_{x}={\bar{Q}}_{x}+C^{\mbox{\tiny T}}Q_{y}C$. Under the rank condition (51), according to Theorem 2, the iterative learning equation (48) with the utility function (69) has a unique solution at any iteration $j$. With the solution to (48), the value evaluation (49) can be rewritten as
$\mathcal{P}_{j+1}=M^{\mbox{\tiny T}}\!\left(A^{\mbox{\tiny T}}P_{j}A\!+\!{\tilde{Q}}_{x}\!-\!A^{\mbox{\tiny T}}P_{j}B\left(R\!+\!B^{\mbox{\tiny T}}P_{j}B\right)^{-1}\!B^{\mbox{\tiny T}}P_{j}A\right)\!M\!+\!{\bar{Q}}_{\varepsilon}.$ (70)
Under the rank condition (63), i.e., $M$ is of full row rank, it follows from (70) that
$(MM^{\mbox{\tiny T}})^{-1}M\mathcal{P}_{j+1}M^{\mbox{\tiny T}}(MM^{\mbox{\tiny T}})^{-1}$
$=$ $A^{\mbox{\tiny T}}P_{j}A\!+\!{\tilde{Q}}_{x}\!-\!A^{\mbox{\tiny T}}P_{j}B\left(R\!+\!B^{\mbox{\tiny T}}P_{j}B\right)^{-1}B^{\mbox{\tiny T}}P_{j}A:=P_{j+1}.$ (71)
The control policy update law at iteration $j$ is given as
${K}_{j}=\left(R\!+\!B^{\mbox{\tiny T}}P_{j}B\right)^{-1}B^{\mbox{\tiny T}}P_{j}A.$ (72)
Substituting (72) into (4.2), one can obtain
${P}_{j+1}-P_{j}=\left(A\!-\!BK_{j}\right)^{\mbox{\tiny T}}P_{j}\left(A\!-\!BK_{j}\right)\!-\!P_{j}\!+\!{\tilde{Q}}_{x}\!+\!K_{j}^{\mbox{\tiny T}}RK_{j}.$
According to Dorato & Levis (1971) that $\lim_{j\rightarrow\infty}P_{j}=P^{*}$, we have
$\lim_{j\rightarrow\infty}\!\left(A\!-\!BK_{j}\right)^{\mbox{\tiny T}}\!P_{j}\left(A\!-\!BK_{j}\right)\!-\!P_{j}\!=\!-{\tilde{Q}}_{x}\!-\!K_{j}^{\mbox{\tiny T}}RK_{j}.$ (73)
Note that ${\tilde{Q}}_{x}>0$ since $Q_{\varepsilon}>0$ and $Q_{y}\geq 0$. If (68) is not satisfied for any iteration $j$, it contradicts the fact shown in (73) where ${\tilde{Q}}_{x}$ is positive definite. This completes the proof.
Finally, by Lemmas 2 and 3, we are ready to show the stability of the system as follows.
Theorem 6
Under the rank condition (51) and (63), there exists an iteration $j$ such that the closed-loop system under the output feedback control $u(k)=-\mathcal{K}_{j}\eta(k)$ is asymptotically stable.
Proof. Under the rank condition (51) and (63), by Lemma 3, for any symmetric positive definite matrix $Q_{\epsilon}$, there always exists an iteration such that (68) holds. Then, there exists a positive definite symmetric matrix $\mathcal{Q}$ satisfying
$M_{\varepsilon_{x}\leftarrow\eta_{\varepsilon}}^{\mbox{\tiny T}}\left(\left(A\!-\!BK_{j}\right)^{\mbox{\tiny T}}\!P_{j}\!\left(A\!-\!BK_{j}\right)\!-\!P_{j}\right)M_{\varepsilon_{x}\leftarrow\eta_{\varepsilon}}=-\mathcal{Q}.$
By Lemma 2, we have $\sigma(A\!-\!BK_{j})\in\mathbb{U}^{-}$. Combining Theorem 1, the closed-loop system under the output feedback control $u(k)=-\mathcal{K}_{j}\eta(k)$ is asymptotically stable. This completes the proof.
Furthermore, combining Lemma 3 with the model-free VI scheme and model-free PI scheme, we propose an output feedback learning control approach with a model-free switched iteration (SI) scheme in Algorithm 5.
Remark 5
Compared to the VI scheme, the PI scheme has a faster convergence rate while the initial admissible control policy is required. The proposed model-free SI scheme provides a way to combine the VI scheme with the PI scheme, which takes advantage of both PI and VI schemes to some extent while avoiding their disadvantages.
Algorithm 5. Model-Free SI-Based Learning Control
Initialization. Define a utility function as $\mathcal{C}(k)$ in (2) and a modified utility function as $\tilde{\mathcal{C}}(k)$ in (69) with $Q_{\varepsilon}=Q_{\varepsilon}^{\mbox{\tiny T}}>0$. Select an pair $(\mathcal{G}_{1},\mathcal{G}_{2})$, an initial internal state $\eta(0)$, an initial control gain $\mathcal{K}_{0}$, and a small threshold $\epsilon>0$. Apply the initial control policy $u_{0}=-{\mathcal{K}_{0}}{\eta}+\xi$ with exploration noise $\xi$. Set $\textbf{sign}_{\rm{PI}}=0$ and $j=0$.
Data Collection. Collect the sampled data $(u(k),y(k))$ and calculate $(\eta(k),\eta(k+1))$ until the rank condition (51) holds.
loop:
  if $\textbf{sign}_{\rm{PI}}=1$
    Calculate $\mathcal{P}_{j}$, $\mathcal{P}_{A,B}^{(j)}$, and $\mathcal{P}_{B,B}^{(j)}$ by (54) with $\mathcal{C}(k)$.
    Update $\mathcal{K}_{j+1}$ by (55).
  else
    Calculate $\mathcal{P}_{\eta,\eta}^{(j)}$, $\mathcal{P}_{\eta,u}^{(j)}$, and $\mathcal{P}_{u,u}^{(j)}$ by (48) with $\tilde{\mathcal{C}}(k)$.
    Update $\mathcal{K}_{j}$ and $\mathcal{P}_{j+1}$ by (50) and (49), respectively.
    if $\mathcal{H}_{j}(P_{j},K_{j})<0$
      $\textbf{sign}_{\rm{PI}}=1$,    continue
    end if
  end if
  if $j\geq 1$ and $\|\mathcal{P}_{j}-\mathcal{P}_{j-1}\|<\epsilon$
    return $\tilde{\mathcal{K}}^{*}=\mathcal{K}_{j}$
    break
  end if
    $j=j+1$
end loop
4.3 Optimality Analysis
To analyze the optimality of several output feedback controllers with respect to the pre-defined value function (2), the optimality analysis among the proposed state parameterization based dynamic output feedback controller (37), the state reconstruction based dynamic output feedback controller (9), the Luenburger observer state parameterization based dynamic output feedback controller (12), and the static output feedback controller is given in this subsection.
According to (2.2) and Theorem 1, note that both the optimal controllers (9) and (37) are equivalent to the following full-state feedback controller
$u_{x}(k)=-{K^{*}}x(k)$ (74)
where $K^{*}$ is obtained by (5). The controller (12) equals to a traditional Luenburger state observer based output feedback controller
$u_{\hat{x}}(k)=-K^{*}\hat{x}(k)$ (75)
where $\hat{x}(k)$ is the estimated state by the Luenburger state observer (10). This implies that the optimality analysis among controllers (9), (12), and (37) can be given by the analysis between the full-state feedback controller (74) and the Luenburger state observer based output feedback controller (75).
Proposition 1
Given any weight matrices $Q_{y}=Q_{y}^{\mbox{\tiny T}}\geq 0$ and $R=R^{\mbox{\tiny T}}>0$, the inequality
$V_{u_{\hat{x}}}(x(k_{0}),\varepsilon_{x}(k_{0}))\geq V_{u_{x}}(x(k_{0})),~{}\forall x(k_{0})\neq\mathbf{0},~{}\varepsilon_{x}(k_{0})\neq\mathbf{0}$
holds, where $V_{u_{x}}(x(k_{0}))$ and $V_{u_{\hat{x}}}(x(k_{0}),\varepsilon_{x}(k_{0}))$ denote the value function (3) at an initial time instant $k_{0}$ under the state feedback controller (74) and the output feedback controller (75), respectively.
Proof. The closed-loop system with controller $u_{x}(k)=-K^{*}x(k)$ can be described as
$x(k+1)=(A-B{K^{*}})x(k).$ (76)
The value function in (3) of the closed-loop system (76) can be rewritten as
$V_{u_{x}}(x(k_{0}))=\sum_{k=k_{0}}^{\infty}x^{\mbox{\tiny T}}(k)\left(Q_{x}+({K^{*}})^{\mbox{\tiny T}}R{K^{*}}\right)x(k)=x^{\mbox{\tiny T}}(k_{0})P_{u_{x}}x(k_{0})$
where $P_{u_{x}}$ is the positive semi-definite solution to
$(A\!-\!BK^{*})^{\mbox{\tiny T}}P_{u_{x}}(A\!-\!BK^{*})\!-\!P_{u_{x}}\!+\!C^{\mbox{\tiny T}}Q_{y}C\!+\!({K^{*}})^{\mbox{\tiny T}}R{K^{*}}\!=\!\mathbf{0}.$ (77)
Define $\bar{x}={\mbox{col}}(x,\varepsilon_{x})$. The closed-loop system with controller $u_{\hat{x}}(k)=-K^{*}\hat{x}(k)$ can be described as
${\bar{x}}(k+1)\!=\!\left[\begin{matrix}A-B{K^{*}}&B{K^{*}}\\ 0&A-LC\end{matrix}\right]\bar{x}(k).$ (78)
The utility function (2) at the instant $k$ under control policy $u_{\hat{x}}(k)$ is given as
$\mathcal{C}_{u_{\hat{x}}}(x(k),\varepsilon_{x}(k))$
$=$ $x^{\mbox{\tiny T}}(k)C^{\mbox{\tiny T}}Q_{y}Cx(k)\!+\!\bar{x}^{\mbox{\tiny T}}(k)\left[\begin{matrix}({K^{*}})^{\mbox{\tiny T}}R{K^{*}}&\!-({K^{*}})^{\mbox{\tiny T}}R{K^{*}}\\ -({K^{*}})^{\mbox{\tiny T}}R{K^{*}}&({K^{*}})^{\mbox{\tiny T}}R{K^{*}}\end{matrix}\right]\bar{x}(k)$
$=$ $\bar{x}^{\mbox{\tiny T}}(k)\left[\begin{matrix}C^{\mbox{\tiny T}}Q_{y}C+({K^{*}})^{\mbox{\tiny T}}R{K^{*}}&-({K^{*}})^{\mbox{\tiny T}}R{K^{*}}\\ -({K^{*}})^{\mbox{\tiny T}}R{K^{*}}&({K^{*}})^{\mbox{\tiny T}}R{K^{*}}\end{matrix}\right]\!\bar{x}(k).$
Then, the value function can be described by
$V_{u_{\hat{x}}}(x(k_{0}),\varepsilon_{x}(k_{0}))=\bar{x}^{\mbox{\tiny T}}(k_{0})\left[\begin{matrix}P_{11}&P_{12}\\ P_{21}&P_{22}\end{matrix}\right]\bar{x}(k_{0}):=\bar{x}^{\mbox{\tiny T}}(k_{0})P_{\hat{u}_{x}}\bar{x}(k_{0}).$
Thus, the Bellman equation can be described as
$\bar{x}^{\mbox{\tiny T}}(k)P_{\hat{u}}\bar{x}(k)$
$=$ $\bar{x}^{\mbox{\tiny T}}(k\!+\!1)P_{\hat{u}}\bar{x}(k\!+\!1)$
$+\bar{x}^{\mbox{\tiny T}}(k)\left[\begin{matrix}C^{\mbox{\tiny T}}Q_{y}C+({K^{*}})^{\mbox{\tiny T}}R{K^{*}}&-({K^{*}})^{\mbox{\tiny T}}R{K^{*}}\\ -({K^{*}})^{\mbox{\tiny T}}R{K^{*}}&({K^{*}})^{\mbox{\tiny T}}R{K^{*}}\end{matrix}\right]\!\bar{x}(k)$ (79)
Substituting (78) into (79), we have
$(A\!-\!BK^{*})^{\mbox{\tiny T}}P_{11}(A\!-\!BK^{*})\!-\!P_{11}\!=\!-C^{\mbox{\tiny T}}Q_{y}C\!-\!{({K^{*}})}^{\mbox{\tiny T}}{R}{{K^{*}}}$ (80a)
${(A\!-\!B{K^{*}})}^{\mbox{\tiny T}}{P_{11}}BK^{*}\!+\!{(A\!-\!B{K^{*}})}^{\mbox{\tiny T}}P_{12}{(A\!-\!LC)}\!-\!{P_{12}}\!=\!{({K^{*}})}^{\mbox{\tiny T}}{R}{{K^{*}}}$ (80b)
${(A\!-\!LC)}^{\mbox{\tiny T}}{P_{21}}{(A\!-\!B{K^{*}})}\!+\!(B{K^{*}})^{\mbox{\tiny T}}{P_{11}}{(A\!-\!B{K^{*}})}\!-\!{P_{21}}\!=\!{({K^{*}})}^{\mbox{\tiny T}}{R}{{K^{*}}}$ (80c)
$(B{K^{*}})^{\mbox{\tiny T}}{P_{11}}B{K^{*}}\!+\!{(A\!-\!LC)}^{\mbox{\tiny T}}{P_{21}}BK^{*}\!+\!(B{K^{*}})^{\mbox{\tiny T}}{P_{12}}{(A\!-\!LC)}$
$+\!{(A\!-\!LC)}^{\mbox{\tiny T}}P_{22}{(A\!-\!LC)}\!-\!P_{22}=\!-\!{({K^{*}})}^{\mbox{\tiny T}}{R}{{K^{*}}}.$ (80d)
It follows from (77) and (80a) that
$P_{11}=P_{u_{x}}$ (81)
Note that $P_{u_{x}}$ is a unique solution to the discrete algebraic Riccati equation (6). From (5) and (81), we have ${(A\!-\!B{K^{*}})}^{\mbox{\tiny T}}{P_{11}}BK^{*}=(B{K^{*}})^{\mbox{\tiny T}}{P_{11}}{(A\!-\!B{K^{*}})}=\left({K^{*}}\right)^{\mbox{\tiny T}}{R}{{K^{*}}}$. Thus, (80b) and (80c) can be rewritten as
${(A\!-\!B{K^{*}})}^{\mbox{\tiny T}}P_{12}{(A\!-\!LC)}\!-\!{P_{12}}~{}=\mathbf{0}$ (82)
${(A\!-\!LC)}^{\mbox{\tiny T}}{P_{21}}{(A\!-\!B{K^{*}})}-P_{21}=\mathbf{0}.$ (83)
With the operators defined in Subsection 2.1, (82) can be expressed as
$\left((A\!-\!LC)^{\mbox{\tiny T}}\otimes(A\!-\!B{K^{*}})^{\mbox{\tiny T}}-I_{n}\otimes I_{n}\right)\mbox{vec}(P_{12})=\mathbf{0}.$ (84)
Since $\sigma(A\!-\!B{K^{*}})\in\mathbb{U}^{-}$ and $\sigma(A\!-\!LC)\in\mathbb{U}^{-}$, the unique solution to (84) is $P_{12}=\mathbf{0}$. In the same way, we have $P_{21}=\mathbf{0}$. Then, (80) can be rewritten as
${(A\!-\!LC)}^{\mbox{\tiny T}}P_{22}{(A\!-\!LC)}\!-\!P_{22}\!=\!-{({K^{*}})}^{\mbox{\tiny T}}(R\!+\!B^{\mbox{\tiny T}}P_{11}B){{K^{*}}}\!.\!$ (85)
Since $\sigma(A\!-\!LC)\in\mathbb{U}^{-}$, there always exists a unique positive semi-definite solution $P_{22}$ to (85). Thus, by (81), the value function $V_{u_{\hat{x}}}(x(k_{0}),\varepsilon_{x}(k_{0}))$ is given as
$V_{u_{\hat{x}}}(x(k_{0}),\varepsilon_{x}(k_{0}))=$ $x^{\mbox{\tiny T}}(k_{0})P_{11}x(k_{0})\!+\!\varepsilon_{x}^{\mbox{\tiny T}}(k_{0})P_{22}\varepsilon_{x}(k_{0})$
$\geq$ $V_{u_{x}}(x(k_{0})),~{}~{}\forall x(k_{0})\neq\mathbf{0},\varepsilon_{x}(k_{0})\neq\mathbf{0}.$
The proof is thus completed.
Proposition 1 shows that the dynamic output feedback controllers (9) and (37) can solve Problem 1, while the controller (12) is not a solution to Problem 1. The key step of the controller design of (9) and (37) is to establish the equivalent relationship between the dynamic output feedback controllers and the full-state feedback controller. The equivalent relationship guarantees the optimality of the dynamic output feedback controller to solve Problem 1.
Distinct from the dynamic output feedback controller, the static output feedback controller described by
$u_{y}(k)=-K_{y}y(k)$ (86)
where $K_{y}$ is the feedback control gain, only needs to access the output data at the current instant. Compared to the proposed dynamic output feedback controller (37), the static output feedback controller may not have the ability to solve Problem 1. We next give the optimality analysis between the dynamic output feedback controller and the static output feedback controller.
To guarantee the feasibility of the static output feedback to the LQR problem, we assume that there exist real matrices $K_{y}$ and $H$ satisfying (Rosinová et al., 2003)
$K_{y}C-(R+B^{\mbox{\tiny T}}P_{u_{y}}B)^{-1}B^{\mbox{\tiny T}}P_{u_{y}}A=H$ (87)
where $P_{u_{y}}$ is the symmetric positive semi-definite solution to
$\mathbf{0}=$ $A^{\mbox{\tiny T}}P_{u_{y}}A\!-\!P_{u_{y}}\!-\!A^{\mbox{\tiny T}}P_{u_{y}}B\left(R\!+\!B^{\mbox{\tiny T}}P_{u_{y}}B\right)^{\!-\!1}B^{\mbox{\tiny T}}PA\!+\!C^{\mbox{\tiny T}}Q_{y}C$
$\!+\!H^{\mbox{\tiny T}}(R+B^{\mbox{\tiny T}}P_{u_{y}}B)H.$ (88)
The above assumption is a necessary and sufficient condition to ensure the stabilizability of the closed-loop system (1) under the static output feedback controller (86). In what follows, we show that the static output feedback control may also be sub-optimal with respect to the utility function (2).
Proposition 2
Given any weight matrices $Q_{y}=Q_{y}^{\mbox{\tiny T}}\geq 0$ and $R=R^{\mbox{\tiny T}}>0$, the inequality
$V_{u_{y}}(x(k_{0}))\geq V_{u_{x}}(x(k_{0})),~{}~{}~{}~{}~{}\forall x(k_{0})\neq\mathbf{0}$ (89)
holds, where $V_{u_{x}}(x(k_{0}))$ and $V_{u_{y}}(x(k_{0}))$ represent the value function (3) under the state feedback controller (74) and the static output feedback controller (86), respectively.
Proof. Let $K^{\prime}_{y}=(R+B^{\mbox{\tiny T}}P_{u_{y}}B)^{-1}B^{\mbox{\tiny T}}P_{u_{y}}A$ and $A_{y}=A-BK^{\prime}_{y}$. It follows from (4.3) that
$\mathbf{0}\!=\!A_{y}^{\mbox{\tiny T}}P_{u_{y}}A_{y}\!-\!P_{u_{y}}\!+\!{K^{\prime}_{y}}^{\mbox{\tiny T}}RK^{\prime}_{y}\!+\!C^{\mbox{\tiny T}}Q_{y}C+\!H^{\mbox{\tiny T}}(R\!+\!B^{\mbox{\tiny T}}P_{u_{y}}B)H$ (90)
Define $\mathcal{P}=A_{y}^{\mbox{\tiny T}}P_{u_{x}}A_{y}-P_{u_{x}}+{K^{\prime}_{y}}^{\mbox{\tiny T}}RK^{\prime}_{y}$, where $P_{u_{x}}$ is the solution to (77). One can obtain
$\mathcal{P}\!=$ $(A-BK^{*})^{\mbox{\tiny T}}P_{u_{x}}(A-BK^{*})-P_{u_{x}}+(K^{*})^{\mbox{\tiny T}}RK^{*}$
$+\left({K^{*}}-K^{\prime}_{y}\right)^{\mbox{\tiny T}}(R+B^{\mbox{\tiny T}}P_{u_{x}}B)\left({K^{*}}-K^{\prime}_{y}\right)$
$=$ $-\!C^{\mbox{\tiny T}}Q_{y}C\!+\!\left({K^{*}}\!-\!K^{\prime}_{y}\right)^{\mbox{\tiny T}}(R\!+\!B^{\mbox{\tiny T}}P_{u_{x}}B)\left({K^{*}}\!-\!K^{\prime}_{y}\right)$ (91)
Subtracting (90) from (4.3), we have
$A_{y}^{\mbox{\tiny T}}(P_{u_{y}}-P_{u_{x}})A_{y}-(P_{u_{y}}-P_{u_{x}})$ (92)
$=$ $-\!H^{\mbox{\tiny T}}(R\!+\!B^{\mbox{\tiny T}}P_{u_{y}}B)H\!-\!\left({K^{*}}\!-\!K^{\prime}_{y}\right)^{\mbox{\tiny T}}(R\!+\!B^{\mbox{\tiny T}}P_{u_{x}}B)\left({K^{*}}\!-\!K^{\prime}_{y}\right).$
Since $\sigma(A_{y})\in\mathbb{U}^{-}$, $R>0$, $P_{u_{y}}\geq 0$, and $P_{u_{x}}\geq 0$, the solution $(P_{u_{y}}-P_{u_{x}})$ to (92) has the property that $(P_{u_{y}}-P_{u_{x}})\geq 0$. Thus, we have
$V_{u_{y}}(x(k_{0}))=x^{\mbox{\tiny T}}(k_{0})P_{u_{y}}x(k_{0})\geq V_{u_{x}}(x(k_{0}))=x^{\mbox{\tiny T}}(k_{0})P_{u_{x}}x(k_{0})$
This completes the proof.
Finally, we give a generalized dynamic output feedback learning control solution to Problem 1 of which the convergence, stability, and optimality can be guaranteed without any prior knowledge of system matrices.
Theorem 7
Under Assumption 1, if the rank condition (51) is satisfied, the dynamic output feedback controller (13), where the control gain is designed by the learning Algorithm 3 or 4 or 5, can solve Problem 1 with guaranteed convergence, stability, and optimality performance.
Proof. Under Assumption 1, the internal model pair $(\mathcal{G}_{1},\mathcal{G}_{2})$ can be user-defined and $\mathcal{G}_{1}$ can be designed to have $n$ distinct eigenvalues, which implies that $A_{L}$ has $n$ distinct eigenvalues. Combining Lemma 1 and the definition of $M$ in (33), it is easy to get that the rank condition (63) holds when $A_{L}$ has $n$ distinct eigenvalues. If the rank conditions (51) and (63) are satisfied, it follows from Theorems 4 and 5 that the estimated control gain $\mathcal{K}_{j}$ by Algorithm 3 or 4 converges to the optimal value $\mathcal{K}^{*}$. The stability performance of Algorithm 4 holds during the whole learning process due to the inherent property of the PI scheme as shown in Hewer (1971). Combining Lemma 3 and Theorems 4 and 5, the convergence and stability performance of Algorithm 5 can be obtained. Moreover, according to Propositions 1 and 2, the control policy (13) where the control gain is estimated by Algorithm 3 or 4 or 5 can solve Problem 1 with optimality. The proof is thus completed.
5 Simulation
In this section, two numerical simulations are given to verify the effectiveness of the proposed output feedback learning control method.
5.1 Unstable and uncontrollable system
Consider the system described by
$x(k+1)=$ $\left[\begin{matrix}1&0.5\\ 0&0.6\end{matrix}\right]x(k)+\left[\begin{matrix}1\\ 0\end{matrix}\right]u(k)$
$y(k)=$ $\left[\begin{matrix}1&1\end{matrix}\right]x(k).$
The initial state is $x(0)={\mbox{col}}\left(0,10\right)$. The eigenvalues of the open-loop system are $1$ and $0.6$, and the system is uncontrollable.
It is shown in Rizvi & Lin (2023) that the convergence of the learning algorithms in Rizvi & Lin (2019) and Rizvi & Lin (2023) cannot be guaranteed when one of the eigenvalues of the observer matrix $A_{L}$ is equal to $0.6$.
To illustrate the effectiveness of our proposed learning algorithm, we chose the eigenvalues of $A_{L}$ as $0.6$ and $0.95$ and chose the initial state of the internal model as $\eta(0)={\mbox{col}}\left(0,0,0,0,5,-5\right)$, which gives the pair $(\mathcal{G}_{1},\mathcal{G}_{2})$ defined in (36) as $\mathcal{G}_{1}=\mbox{block diag}[\mathcal{A},\mathcal{A},\mathcal{A}_{\varepsilon}],~{}\mathcal{G}_{2}=\mbox{block diag}[b,b,\mathbf{0}]$ with
$\mathcal{A}_{\varepsilon}\!=\!\mathcal{A}\!=\!\left[\begin{array}[]{ccc}0&1\\ -0.57&1.55\end{array}\right],~{}~{}b\!=\!\left[\begin{array}[]{ccc}0\\ 1\end{array}\right]$
and the state parameterization matrices
$M_{u}=\left[\begin{matrix}-0.6&1\\ 0&0\end{matrix}\right],M_{y}=\left[\begin{matrix}-0.03&0.05\\ 0&0\end{matrix}\right]$
$M_{\varepsilon_{x}\leftarrow\eta_{\varepsilon}}=\left[\begin{matrix}-0.2885&-0.2885\\ 0.9744&-1.0256\end{matrix}\right].$
Then, we set the weight matrices $Q_{y}=1$ and $R=1$, which gives the optimal control gain ${\mathcal{K}}^{*}$ of dynamic output feedback controller computed as
${\mathcal{K}}^{*}\!=\![-0.3708,~{}0.6180,-0.0185,~{}0.0309,~{}0.5020,-0.8944]$
and $\|\mathcal{P}^{*}\|=6.1046$.
The parameters in Algorithm 3 are chosen as $\mathcal{P}_{0}=I_{n_{z}}$ with $n_{z}=6$, $\mathcal{K}_{0}=\mathbf{0}$, and $\epsilon=0.001$. The exploration noise is a combination of $100$ sinusoids with random frequency $w_{i}$, amplitude $c_{i}$, and initial phase $d_{i}$, i.e., $\xi=\sum_{i=1}^{100}c_{i}\mbox{sin}(w_{i}k+d_{i})$. Two numerical simulations are carried out by employing Algorithm 3 with the proposed dynamic output feedback controller (denoted as $u$) and the traditional Luenburger observer-based output feedback controller (denoted as $\hat{u}$). The simulation results are shown in Figs. 33. Fig. 3 depicts the convergence performance of Algorithm 3, which shows that, after $83$ iterations, the estimated control gain is updated as
${\tilde{\mathcal{K}}}^{*}=[-0.3708,~{}0.6181,~{}-0.0186,~{}0.0309,~{}0.5021,~{}-0.8945].$
The evolution of inputs and outputs under the output feedback controllers $u$ and $\hat{u}$ is drawn in Fig. 3, which shows that the controllers are updated at $k_{N}=29$ when the rank condition is met. The evolution of the observer error is shown in Fig. 3. It follows from Figs. 3 and 3 that the transient performance of the closed-loop system under controller $u$ is better than that of the closed-loop system under controller $\hat{u}$.
Fig. 1: The convergence of Algorithm 3.
Fig. 2: The trajectories of inputs and outputs under different controllers.
Fig. 3: The evolution of observer error.
5.2 Application to the aircraft system
Consider the aircraft system dynamics (Rizvi & Lin, 2019) linearized as
$\left[\begin{matrix}\alpha(k\!+\!1)\\ q(k\!+\!1)\\ \delta_{e}(k\!+\!1)\end{matrix}\right]\!=$ $\left[\begin{matrix}0.906488&0.0816012&-0.0005\\ 0.0741349&0.90121&-0.0007083\\ 0&0&0.132655\end{matrix}\right]\left[\begin{matrix}\alpha(k)\\ q(k)\\ \delta_{e}(k)\end{matrix}\right]$
$\!+\!\left[\begin{matrix}-0.00150808\\ -0.0096\\ 0.867345\end{matrix}\right]\!u(k)$
where $\alpha$, $q$, and ${\delta}_{e}$ denote the angle of attack, the pitch rate, and the elevator deflection angle, respectively. $u$ is the control input. $y=\alpha$ is the output. The initial state is ${\mbox{col}}(\alpha(0),q(0),\delta_{e}(0))={\mbox{col}}\left(0.2,0.2,0.2\right)$.
Now, we choose the eigenvalues of the observer matrix $A_{L}$ as $-0.91$, $-0.92$, and $-0.93$, and select the initial state of internal model as $\eta(0)={\mbox{col}}\left(0,0,0,0,0,0,2,1,1\right)$, which gives the pair $(\mathcal{G}_{1},\mathcal{G}_{2})$ defined in (36) as $\mathcal{G}_{1}=\mbox{block diag}[\mathcal{A},\mathcal{A},\mathcal{A}_{\varepsilon}],~{}\mathcal{G}_{2}=\mbox{block diag}[b,b,\mathbf{0}]$ with
$\mathcal{A}_{\varepsilon}\!=\!\mathcal{A}\!=\!\left[\!\begin{array}[]{ccc}0&1&0\\ 0&0&1\\ -0.7786&-2.5391&-2.7600\end{array}\!\right],~{}~{}b\!=\!\left[\!\begin{array}[]{ccc}0\\ 0\\ 1\end{array}\!\right]$
and the state parameterization matrices
$M_{u}$ $\!=\!\left[\begin{array}[]{ccc}0.0003&0.0003&-0.0015\\ 0.0391&0.1145&-0.0096\\ 2.0276&7.8964&0.8673\end{array}\right]$
$M_{y}$ $\!=\!\left[\begin{array}[]{ccc}0.8862&1.4884&4.7004\\ 14.1018&-105.7656&99.6818\\ 2896.8322&-6457.8672&3572.4259\end{array}\right]$
$M_{\varepsilon_{x}\leftarrow\eta_{\varepsilon}}$ $\!=\!\left[\begin{array}[]{ccc}0.2505&-0.4134&0.1123\\ 14.1130&-14.2814&-13.7445\\ 858.3455&-505.9328&-1210.5582\end{array}\right].$
Fig. 4: The convergence of VI-based Algorithm 3.
Fig. 5: The convergence of PI-based Algorithm 4.
Fig. 6: The convergence of SI-based Algorithm 5.
Then, we set the weight matrices $Q_{y}=100$ and $R=1$, which gives the optimal control gain ${\mathcal{K}}^{*}$ of dynamic output feedback controller computed as
${\mathcal{K}}^{*}=[$ $-0.1029,~{}-0.2900,~{}0.0377,~{}-30.0765,~{}264.5137,$
[…]
and $\|\mathcal{P}^{*}\|=4.2444\times 10^{6}$.
Fig. 7: The evolution of input and output.
Fig. 8: The evolution of states under different controllers.
Fig. 9: The evolution of the observer error.
To illustrate the effectiveness of our proposed method, we made three numerical simulations by employing Algorithms 3, 4, and 5, respectively.
The parameters are chosen as $\mathcal{P}_{0}=10^{5}I_{n_{z}}$ with $n_{z}=9$, and $\epsilon=1$. The exploration noise is a combination of $100$ sinusoids with random frequency $w_{i}$, amplitude $c_{i}$, and initial phase $d_{i}$, i.e., $\xi=\sum_{i=1}^{100}c_{i}\mbox{sin}(w_{i}k+d_{i})$. The admissible initial control gain is set as $\mathcal{K}_{0}=\mathbf{0}$ in Algorithm 4, and we choose $\mathcal{K}_{0}^{{\mbox{\tiny T}}}={\mbox{col}}(1.1069,~{}3.3440,~{}-0.1596,~{}469.8189,~{}-3114.9114,~{}2609.3162,~{}399.8526,~{}-382.6236,~{}-418.7970)$ in Algorithms 3 and 5 which is not a stabilizing control gain. It should be noted that we set $Q_{\varepsilon}=0.01I_{n_{z}}$ in Algorithm 5. The simulation results are shown in Figs. 69. Fig. 6 depicts the convergence performance of VI-based Algorithm 3, which shows that, after $241$ iterations, the estimated control gain is updated as
${\tilde{\mathcal{K}}}^{*}=[$ $-0.1033,~{}-0.2910,~{}0.0371,~{}-30.0673,~{}264.4322,~{}$
[…]
It is shown in Fig. 6 that, for PI-based Algorithm 4, the estimated control gain is updated after only $7$ iterations, and its value is
${\tilde{\mathcal{K}}}^{*}=[$ $-0.1027,~{}-0.2896,~{}0.0380,~{}-30.0794,~{}264.5498,~{}$
[…]
The convergence performance of SI-based Algorithm 5 is drawn in Fig. 6, which shows that, after $99$ iterations, the estimated control gain is updated as
${\tilde{\mathcal{K}}}^{*}=[$ $-0.1034,~{}-0.2913,~{}0.0369,~{}-30.0650,~{}264.4091,~{}$
[…]
The main differences among Algorithms 3, 4, and 5 are twofold: 1) the convergence rate; and 2) the requirement of an initial admissible control policy. The above three numerical simulation results show that the SI-based Algorithm 5 is faster than the VI-based Algorithm 3, and it does not need the initial admissible control policy as the PI-based Algorithm 4 does.
In addition, similar to Subsection 5.1, two numerical simulations are carried out by employing Algorithm 3 with $u$ and $\hat{u}$. The evolution of inputs and outputs of the closed-loop system is drawn in Fig. 9, which shows that the controllers are updated at $k_{N}=58$ when the rank condition is met. The trajectory of states under the proposed dynamic output feedback controller $u$ and the traditional Luenburger observer-based output feedback controller $\hat{u}$ are shown in Fig 9. The evolution of observer error during time instant $k\in[150,400]$ is shown in Fig. 9. From Figs. 99, we can see that the proposed output feedback learning control approach is immune to the observer error, while the transient performance of the closed-loop systems under the Luenburger observer-based output feedback controller is influenced by the observer error.
6 Conclusion
This paper presents the optimal solution to the discrete-time LQR problem with unknown system matrices and unmeasurable states. The significance of this result is that a generalized dynamic output feedback learning control method is proposed. The model-free learning algorithms are established by employing the PI scheme, the VI scheme, and the SI scheme to estimate the optimal control policy to solve the LQR problem. In addition, we conducted in-depth discussions regarding the theoretical guarantee of the proposed ADP-based output feedback learning control methods, including convergence, stability, and optimality.